### v1

# Ordinary primes for $\operatorname{GL}_{2}$-type abelian varieties and weight $2$ modular forms

Abstract.
Let $A$ be a $g$-dimensional abelian variety defined over a number field $F$. It is conjectured that the set of ordinary primes of $A$ over $F$ has positive density, and this is known to be true when $g=1,2$, or for certain abelian varieties with extra endomorphisms. In this paper, we extend the family of abelian varieties whose sets of ordinary primes have positive density. Specifically, we show that if the endomorphism algebra of $A$ contains a number field $K$ of degree $g$, then under certain conditions on the fields $F$ and $K$, the set of ordinary primes of $A$ over $F$ has positive density. This includes $\operatorname{GL}_{2}$-type abelian varieties over $\mathbb{Q}$ (resp. quadratic number fields) of dimension $q$ or $2q$ (resp. $q$) for any rational prime $q$. The proof is carried out in the general setting of compatible systems of Galois representations, and as a consequence, it also implies a positive density result for the sets of ordinary primes of certain modular forms of weight $2$.
2010 Mathematics Subject Classification:
Primary 11G10, 11R45, 11F80, 11F30; Secondary 14K15.
1. Introduction
Let $A$ be an abelian variety of dimension $g$ over a number field $F$. Let $v$ be a prime of $F$ that is also a prime of good reduction for $A$. We denote by $A_{v}$ the reduction of $A$ at $v$ and by $P_{A,v}(X)$ the characteristic polynomial of the Frobenius endomorphism of $A_{v}$. Then, $P_{A,v}(X)$ is a polynomial of degree $2g$ with integer coefficients. We write the coefficient of the middle term $X^{g}$ in $P_{A,v}(X)$ as $a_{v,g}(A)$. Then, the prime $v$ is an ordinary prime of $A$ over $F$ if $p\nmid a_{v,g}(A)$, where $p$ is the rational prime below $v$. In this case, $A_{v}$ is an ordinary abelian variety over $\mathbb{F}_{v}$ and we say $A$ is ordinary at $v$. For example, if $A$ is an elliptic curve over $\mathbb{Q}$, then a rational prime $v=p$ is an ordinary prime of $A$ if and only if $p\nmid a_{p,1}(A)$, where $a_{p,1}(A)=p+1-|A_{p}(\mathbb{F}_{p})|$ is the Frobenius trace of $A$ at $p$. When $p\geq 5$, this is also equivalent to that $a_{p,1}(A)\neq 0$ by the Hasse–Weil bound.
The study of ordinary primes for abelian varieties plays a key role in arithmetic geometry. For example, ordinary abelian varieties over $\overline{\mathbb{F}_{v}}$ are generic in the moduli space of all principally polarized abelian varieties [NO80], they exhibit canonical lifting properties to characteristic 0 via Serre–Tate theory [Kat81], and they are crucial in the proof of Hecke orbit conjecture for certain Shimura varieties [Cha05, Cha06].
A comparable definition of “ordinary” also appears in the study of modular forms. Let $f$ be a newform of weight $k\geq 2$ and level $N$, let $a_{n}(f)$ be the $n$-th Fourier coefficient of $f$, and let $K_{f}$ be the coefficient field of $f$. For a rational prime $p\nmid N$ and a prime $\lambda|p$ of $K_{f}$, $f$ is called $\lambda$-ordinary if $\lambda\nmid(a_{p}(f))$, and $f$ is called $p$-ordinary if $f$ is $\lambda$-ordinary for all $\lambda|p$, or equivalently, $\lambda\nmid(a_{p}(f))$ for all primes $\lambda|p$ of $K_{f}$. We will use the phrases “$f$ is $p$-ordinary”, “$f$ is ordinary at $p$”, and “$p$ is an ordinary prime of $f$ over $\mathbb{Q}$” interchangeably.
The notion of ordinary modular forms first appeared in a paper by Hida [Hid81]. This paper was followed by a series work of Hida, where he established the theory of ordinary $p$-adic modular forms, or nowadays known as Hida theory. The ordinary assumption has since played an important role in the theory of ($p$-adic) modular forms, particularly in Iwasawa theory, e.g., in Emerton–Pollack–Weston [EPW06] and Skinner–Urban [SU14] (and many others). The definition of ordinariness that we adopt in this paper is, however, different from (but closely related to) Hida’s definition. Roughly speaking, Hida’s definition of $p$-ordinariness only requires $a_{p}(f)$ to be a $p$-adic unit under a prechosen embedding $K_{f}\hookrightarrow\overline{\mathbb{Q}_{p}}$, while our definition requires $a_{p}(f)$ to be a $p$-adic unit under all embeddings $K_{f}\hookrightarrow\overline{\mathbb{Q}_{p}}$. In particular, our definition of $p$-ordinariness coincides with the one by Hida if $K_{f}=\mathbb{Q}$. The reason for us to adopt a different definition is that a weight $2$ newform $f$ is ordinary at $p$ under our definition if and only if the abelian variety $A_{f}/\mathbb{Q}$ associated to $f$ via the Eichler–Shimura theory is ordinary at $p$ (see Proposition 3.12).
In this paper, we prove a positive density result for the sets of ordinary primes of certain abelian varieties and modular forms by analyzing under the framework of Galois representations.
We begin by reviewing prior works on the study of ordinary primes for an abelian variety $A$ over a number field $F$. A conjecture commonly attributed to Serre [DMOS82, Proposition 2.6, p. 370] predicts that there are infinitely many ordinary primes for $A/F$. Furthermore, it is conjectured that the density of ordinary primes for $A$ over $\overline{F}$ is 1. The conjecture is proved when the dimension of $A$ is 1 or 2 [Ser00, DMOS82, Saw16]. However, much less is known for higher dimensional abelian varieties. Nonetheless, the conjecture holds for special cases where $A$ is not generic, such as the case when $A/F$ is a CM abelian variety [Fit24, Remark 12] (i.e., the geometric endomorphism algebra $\operatorname{End}_{\overline{F}}(A)\otimes\mathbb{Q}$ is a CM field of degree $2g$) or when the Mumford–Tate group of $A$ is small in the sense of [Noo95, Theorem 2.2, p. 168] or [Pin98, Theorem 7.1].
A recent work by Fité [Fit24, Theorem 3] showed that the sets of ordinary primes of certain modular abelian varieties have positive density. More generally, Fité [Fit24, Theorem 1, Theorem 4, and Corollary 5] proved this result for abelian varieties that have additional endomorphisms. In comparison with Fité’s work on modular abelian varieties, the work by Suh [Suh20, Theorem, p. 649] showed that the sets of ordinary primes of certain Hilbert newforms of parallel weight 2 have positive density. The reason for this comparison is that Hilbert newforms of parallel weight $2$ are conjecturally associated to certain abelian varieties over totally real number fields with extra endomorphisms. More recently, the forthcoming work by Cantoral Farfán, Li, Mantovan, Pries, and Tang established a positive density result for the sets of ordinary primes of certain abelian varieties parametrized by unitary Shimura curves with simple signature.
We extend the family of higher dimensional abelian varieties whose sets of ordinary primes have positive density. Our first two results are for $\operatorname{GL}_{2}$-type abelian varieties over $\mathbb{Q}$ and quadratic number fields of specific dimensions.
Theorem 1.1.
Let $A/\mathbb{Q}$ be an abelian variety of dimension $q$ (resp. $2q$) for some rational prime $q$. Assume that $\operatorname{End}_{\mathbb{Q}}(A)\otimes\mathbb{Q}$ contains a number field $K$ of degree $q$ (resp. $2q$). Then, the set of ordinary primes of $A$ over $\mathbb{Q}$ has positive density.
Theorem 1.2.
Let $A/F$ be an abelian variety of dimension $q$ over a quadratic number field $F$ for some rational prime $q$. Assume that $\operatorname{End}_{F}(A)\otimes\mathbb{Q}$ contains a number field $K$ of degree $q$. Then, the set of ordinary primes of $A$ over $F$ has positive density.
More generally, let $A/F$ be a simple abelian variety of an arbitrary dimension $g$ with whose endomorphism algebra contains a number field $K$ of degree $g$. The second result gives an affirmative answer to the positive density conjecture on ordinary primes for $A$ over $F$ under certain restrictions on $F$ and $K$.
Theorem 1.3.
Let $A/F$ be an abelian variety of dimension $g$ over a number field $F$. Assume that
  1. (1)
    $A$ is absolutely simple, i.e., $A/\overline{F}$ is simple;
  2. (2)
    $\operatorname{End}_{F}(A)\otimes\mathbb{Q}$ contains a number field $K$ of degree $g$;
  3. (3)
    $A/\overline{F}$ is of the first kind, i.e., $A/\overline{F}$ is not of type IV in the Albert’s classification.
Suppose further that there exists a rational prime $p$ that lies below a degree $1$ prime of $F$ and either
  1. (P1)
    remains inert in $K$, or
  2. (P2)
    splits into two distinct primes in $K$ of the same inertia degree.
Then, the set of ordinary primes of $A$ over $F$ has positive density.
Remark 1.4.
We provide an example of a family of hyperelliptic curves whose Jacobians satisfy the condition (P1) or (P2). Consequently, the sets of ordinary primes of these abelian varieties have positive density.
The family we consider is taken from [TTV91]. Let $p$ be a rational prime with $p\neq 2$ or $5$. Let $\zeta_{p}$ denote the primitive $p$-th root of unity in $\overline{\mathbb{Q}}$, $g(X)\in\mathbb{Z}[X]$ be the minimal polynomial of $-(\zeta_{p}+\zeta_{p}^{-1})$, and $f_{t}(X):=Xg(X^{2}-2)+t\in\mathbb{Z}[X]$ for any $t\in\overline{\mathbb{Q}}$. The equation
$C_{t}:y^{2}=f_{t}(x)$
characterizes a family of absolutely simple hyperelliptic curves of genus $\frac{p-1}{2}$ [TTV91, Corollary 6]. For each curve $C_{t}$, the Jacobian variety $J_{C_{t}}$ is defined over $F:=\mathbb{Q}(t)$, of dimension $\frac{p-1}{2}$, and contains the totally real field $K:=\mathbb{Q}(\zeta_{p}+\zeta_{p}^{-1})$ in its endomorphism algebra [TTV91, Theorem 1]. Note that $[K:\mathbb{Q}]=\frac{p-1}{2}=\dim J_{C_{t}}$. Since the Galois extension $K/\mathbb{Q}$ is cyclic when $p\geq 3$, if $\ell$ is inert in $K$, then the order of $\ell\pmod{p}$ is divisible by $\frac{p-1}{2}$, and if $\ell$ splits into two primes of the same inertia degree, then the order of $\ell\pmod{p}$ is divisible by $\frac{p-1}{4}$ (if $p\equiv 1\pmod{4}$). Clearly, such a prime $\ell$ exists. Thus, the condition (P1) is satisfied if $p\geq 3$, and the condition (P2) is satisfied if $p\equiv 1\pmod{4}$. Finally, by choosing a number field $F$ such that $\ell$ lies below a degree $1$ prime of $F$ and taking $t$ as the primitive element of $F$, we see that $\{J_{C_{t}}\}_{t\in F}$ is a family of abelian varieties whose sets of ordinary primes over $F$ have positive density.
In the case of modular forms, we obtain the following result under our definition of ordinary primes of modular forms.
Theorem 1.5.
Let $f$ be a newform of weight $2$ and level $N$ and let $K_{f}$ denote the coefficient field of $f$. Suppose that $[K_{f}:\mathbb{Q}]=q$ or $2q$ for some prime $q$. Then, the set of ordinary primes of $f$ over $\mathbb{Q}$ has positive density.
More generally, if there exists a rational prime $p$ that remains inert in $K_{f}$ or splits into two distinct primes in $K_{f}$ of the same inertia degree, then the set of ordinary primes of $f$ over $\mathbb{Q}$ has positive density.
To prove Theorem 1.1, Theorem 1.5 in the case of “$2q$” and Theorem 1.3 in the case of (P2), we begin by proving Theorem 1.6, a general density result for compatible systems of Galois representations. We then apply this theorem to the $\ell$-adic Galois representations of abelian varieties and modular forms.
Theorem 1.6.
Let $F$ and $K$ be two number fields, let $S$ be a finite set of primes of $F$, and let $(a_{v})_{v\in\Sigma_{F}\setminus S}$ be a fixed sequence in $\mathcal{O}_{K}$. Let $\epsilon$ be a finite order Hecke character of $F$ with image lying in $K^{\times}$. Let $T$ be an infinite set of rational primes which split completely in $K$. For each rational prime $\ell\in T$, let $\rho_{\ell}:\operatorname{Gal}(\overline{F}/F)\rightarrow\operatorname{GL}_{2}(K\otimes\mathbb{Q}_{\ell})$ be a continuous Galois representation. Suppose that
  1. (1)
    (ramification) for every $\ell\in T$, each $\rho_{\ell}$ is unramified at every $v\in\Sigma_{F}$ with $v\notin S$ and $v\nmid\ell$;
  2. (2)
    ($\ell$-independence) for every $\ell\in T$ and every $v\in\Sigma_{F}$ with $v\notin S$ and $v\nmid\ell$, the characteristic polynomial of $\rho_{\ell}(\operatorname{Frob}_{v})\in\operatorname{GL}_{2}(K\otimes\mathbb{Q}_{\ell})$ is
    $X^{2}-(a_{v}\otimes 1)X+(\epsilon(v)\mathrm{N}(v)\otimes 1),$
    which is independent of $\ell$, where $\mathrm{N}(v)=|\mathcal{O}_{F}/v|$;
  3. (3)
    (Ramanujan’s conjecture/Hasse–Weil bound) the sequence $(a_{v})_{v\in\Sigma_{F}\setminus S}$ satisfies that
    $|\iota(a_{v})|_{\mathbb{C}}\leq 2\mathrm{N}(v)^{1/2}$
    for all $v\in\Sigma_{F}$ with $v\notin S$ and all embeddings $\iota:K\hookrightarrow\mathbb{C}$;
  4. (4)
    (surjectivity) there exists two finite extensions of fields $F^{\prime}/F$ and $K/K^{\prime}$ such that for every $\ell\in T$, we have, up to conjugation, that
    $\rho_{\ell}(\operatorname{Gal}(\overline{F}/F^{\prime}))=\{M\in\operatorname{GL}_{2}(\mathcal{O}_{K^{\prime}}\otimes\mathbb{Z}_{\ell}):\operatorname{det}M\in(\mathbb{Z}\otimes\mathbb{Z}_{\ell})^{\times}(\simeq\mathbb{Z}_{\ell}^{\times})\}.$
Suppose further that there exists a rational prime $p$ that lies below a degree $1$ prime of $F$ and splits into two distinct primes in $K$ of the same inertia degree. Then, the set
$S^{\mathrm{ord}}:=\{v\in\Sigma_{F}:v\notin S\text{ and }\lambda\nmid(a_{v})\text{ for all $\lambda\in\Sigma_{K}$ with }\lambda|\mathrm{N}(v)\}$
has positive density.
The general approach to proving such results can be traced back to the work of Serre, Katz, and Ogus, and has been successfully adapted with modifications by Pink, Noot, and others, as previously mentioned. Our proof of Theorem 1.6 is also similar and we sketch it as follows. First, we demonstrate that there is a positive density set $S_{0}$ of degree 1 primes $v$ of $F$, lying above rational primes $p$, that split into two distinct primes in $K$ of the same inertia degree. This requires analysis of the splitting behaviors of primes in extensions of number fields (that are not necessarily Galois). Next, we associate to each prime $v\in S_{0}$ the value $\mathrm{N}_{K/\mathbb{Q}}(a_{v})^{2}\cdot\mathrm{N}(v)^{-g}$ and prove that if $v\notin S^{\mathrm{ord}}$, then $\mathrm{N}_{K/\mathbb{Q}}(a_{v})^{2}\cdot\mathrm{N}(v)^{-g}$ belongs to a finite set (with size at most $12^{2g}+$ 1). The proof of finiteness relies on the splitting behavior of $p$ and the bound bound on the archimedean norm of $a_{v}$. Finally, we perform explicit computations using the images of residual Galois representations and the Chebotarev density theorem to show that the density of the primes $v\in S_{0}$ with $v\notin S^{\mathrm{ord}}$ is 0. This step involves studying the images of the $\lambda$-adic Galois representations of abelian varieties and modular forms.
Though Theorem 1.1, Theorem 1.3, and Theorem 1.5 are all applications of Theorem 1.6, we will present a simpler proof of Theorem 1.1 by using the modularity of $\operatorname{GL}_{2}$-type abelian varieties over $\mathbb{Q}$ and Theorem 1.5. Additionally, we emphasize that the surjectivity condition (4) in Theorem 1.6 is intentionally relaxed to avoid the need for assuming that the endomorphisms of $A$ are defined over the base field and that the modular forms have no inner twists.
On the other hand, the proofs of Theorem 1.1, Theorem 1.2, and Theorem 1.5 in the case of “$q$”, and the proof of Theorem 1.3 in the case of (P1) are much simpler. In this case, we only need the following density result for a sequence of algebraic integers indexed by primes.
Theorem 1.7.
Let $F$ and $K$ be two number fields and let $S$ be a finite set of primes of $F$. Let $(a_{v})_{v\in\Sigma_{F}\setminus S}$ be a fixed sequence in $\mathcal{O}_{K}$ satisfying that
  1. (1)
    (Ramanujan’s conjecture) $|\iota(a_{v})|_{\mathbb{C}}\leq 2\mathrm{N}(v)^{1/2}$ for all $v\in\Sigma_{F}$ with $v\notin S$ and all embeddings $\iota:K\hookrightarrow\mathbb{C}$;
  2. (2)
    (non-lacunary) the set $\{v\in\Sigma_{F}:v\notin S\text{ and }a_{v}=0\}$ of primes of $F$ has density 00.
Suppose further that there exists a rational prime $p$ that lies below a degree $1$ prime of $F$ and remains inert in $K$. Then, the set
$S^{\mathrm{ord}}:=\{v\in\Sigma_{F}:v\notin S\text{ and }\lambda\nmid(a_{v})\text{ for all $\lambda\in\Sigma_{K}$ with }\lambda|\mathrm{N}(v)\}$
has positive density.
Remark 1.8.
One can also apply Theorem 1.6 and Theorem 1.7 to obtain similar positive density results on the ordinary primes of certain Hilbert newforms, where one assumes a condition similar to that in Theorem 1.3 on the existence of some rational prime with specific joint splitting behaviors in the base field and in the coefficient field of the Hilbert newform. The main ingredients for applying Theorem 1.6 and Theorem 1.7 to the case of Hilbert modular forms are given by [Car86] and [Tay89] on the existence of the associated Galois representations, [Nek12, Appendix B] on the surjectivity of the Galois images, and [Bla06] on Ramanujan’s conjecture.
Overview
In Section 2, we collect results concerning the splitting types of primes in number field extensions. In Section 3, we review the essential background and results related to the Galois representations of abelian varieties and modular forms. Subsequently, in Section 4 and Section 5, we prove Theorem 1.7 and Theorem 1.6, respectively. The proof of the former relies solely on results from Section 2, while the latter involves also a detailed analysis of the images of the Galois representations. Lastly, in Section 6, we present the proofs of Theorem 1.1, Theorem 1.2, Theorem 1.3, and Theorem 1.5 as applications of Theorem 1.6 and Theorem 1.7. Finally, in the Appendix, we provide an independent proof of the density of ordinary primes (as defined in this paper) for CM modular forms of weight 2.
Notation
Throughout this paper, $A$ usually denotes an abelian variety and $f$ usually denotes a modular form.
In the case of abelian varieties, $g$ usually denotes the dimension of $A$, $F$ usually denotes the field of definition of $A$, and $K$ usually denotes the number field embedded into the $F$-endomorphism algebra of $A$. When enlarging the field of definition of $A$, we usually use $F^{\prime}$ to denote the enlarged field of definition and $K^{\prime}$ to denote the number field embedded into the center of the $F^{\prime}$-endomorphism algebra of $A$ (and usually $K^{\prime}\subseteq K$).
In the case of modular forms, $K_{f}$ usually denotes the coefficient field of $f$ and $g$ usually denotes the degree of $K_{f}$. We usually use the notations $F_{f}^{\prime}/\mathbb{Q}$ and $K_{f}/K_{f}^{\prime}$ for finite extensions of fields, which are needed when discussing the Galois images of the modular forms with inner twists.
For any number field $E$ (e.g., $F,F^{\prime},K,K^{\prime}$), $\mathcal{O}_{E}$ denotes the ring of integers of $E$, $\Sigma_{E}$ denotes the set of nonzero prime ideals of $E$, $\mathrm{N}_{E/\mathbb{Q}}(\cdot)$ denotes the absolute norm of an element of $E$, and $\mathrm{N}(\cdot)$ denotes the ideal norm. Specifically for $F$ (resp. $F^{\prime}$) and $K$ or $K_{f}$ (resp. $K^{\prime}$), we use $v$ (resp. $v^{\prime}$) to denote a prime of $F$ (resp. $F^{\prime}$), use $p$ to denote the rational prime that lies below $v$ (resp. $v^{\prime}$), use $\lambda$ (resp. $\lambda^{\prime}$) to denote a prime of $K$ or $K_{f}$ (resp. $K^{\prime}$), and use $\ell$ to denote the rational prime that lies below $\lambda$ (resp. $\lambda^{\prime}$).
2. Lemmas on primes and number fields
We first begin by proving some lemmas, mainly on the splitting behaviors of primes in (non-Galois) field extensions. In particular, Proposition 2.7 (resp. Proposition 2.9) is crucial to constructing a positive density set of primes that are ordinary in Theorem 1.6 and Theorem 1.7 (resp. Theorem 1.1, Theorem 1.2, and Theorem 1.5).
2.1. Correspondence between places and embeddings
Lemma 2.1.
Let $K/\mathbb{Q}$ be a number field and let $\ell$ be a rational prime. For each embedding $\tau:K\hookrightarrow\overline{\mathbb{Q}_{\ell}}$, define a place $\lambda_{\tau}$ of $K$ by
$|x|_{\lambda_{\tau}}:=|\tau(x)|_{\ell}$
for all $x\in K$, where $|\cdot|_{\ell}$ is the $\ell$-adic absolute value on $\overline{\mathbb{Q}_{\ell}}$ with $|\ell|_{\ell}=\ell^{-1}$. Then, the assignment $\tau\mapsto\lambda_{\tau}$ gives a one-to-one correspondence between embeddings $\tau:K\hookrightarrow\overline{\mathbb{Q}_{\ell}}$ up to pre-composition with elements in $\operatorname{Gal}(\overline{\mathbb{Q}_{\ell}}/\mathbb{Q}_{\ell})$ and places $\lambda|\ell$ of $K$.
Proof.
See (8.1) Extension Theorem in [Neu99, Chapter II]. ∎
Lemma 2.2.
Let $K/\mathbb{Q}$ be a number field and let $\ell$ be a rational prime that splits completely in $K$. Fix an embedding $\iota_{\ell}:\overline{\mathbb{Q}}\hookrightarrow\overline{\mathbb{Q}_{\ell}}$. For each embedding $\sigma:K\hookrightarrow\overline{\mathbb{Q}}$, define a place $\lambda_{\sigma}$ of $K$ by
$|x|_{\lambda_{\sigma}}:=|\iota_{\ell}(\sigma(x))|_{\ell}$
for all $x\in K$. Then, the assignment $\sigma\mapsto\lambda_{\sigma}$ defines a one-to-one correspondence between embeddings $\sigma:K\hookrightarrow\overline{\mathbb{Q}}$ and places $\lambda|\ell$ of $K$.
Proof.
By Lemma 2.1, it suffices to show that the assignment $\sigma\mapsto\iota_{\ell}\circ\sigma$ gives a one-to-one correspondence between embeddings $\sigma:K\hookrightarrow\overline{\mathbb{Q}}$ and embeddings $\tau:K\hookrightarrow\overline{\mathbb{Q}_{\ell}}$ up to pre-composition with elements in $\operatorname{Gal}(\overline{\mathbb{Q}_{\ell}}/\mathbb{Q}_{\ell})$. Let $\sigma:K\hookrightarrow\overline{\mathbb{Q}}$ be an embedding and let $\tau=\iota_{\ell}\circ\sigma$. Since $\ell$ splits completely in $K$, the completion $\widehat{\tau(K)}$ of $\tau(K)$ in $\overline{\mathbb{Q}_{\ell}}$ is thus the same as $\mathbb{Q}_{\ell}$. In this way, $\tau$ factors as
$\tau:K\xrightarrow{\;\sim\;}\tau(K)\subseteq\widehat{\tau(K)}=\mathbb{Q}_{\ell}\subseteq\overline{\mathbb{Q}_{\ell}},$
so $\alpha\circ\tau=\tau$ for any $\alpha\in\operatorname{Gal}(\overline{\mathbb{Q}_{\ell}}/\mathbb{Q}_{\ell})$.
Now, suppose that $\sigma_{1},\sigma_{2}:K\hookrightarrow\overline{\mathbb{Q}}$ satisfy that $\iota_{\ell}\circ\sigma_{1}=\alpha\circ\iota_{\ell}\circ\sigma_{2}$ for some $\alpha\in\operatorname{Gal}(\overline{\mathbb{Q}_{\ell}}/\mathbb{Q}_{\ell})$. Then, $\iota_{\ell}\circ\sigma_{1}=\iota_{\ell}\circ\sigma_{2}$ by the previous discussion, and hence $\sigma_{1}=\sigma_{2}$. This implies that the assignment $\sigma\mapsto\iota_{\ell}\circ\sigma$ is injective, and hence it is bijective by Lemma 2.1 since $\ell$ splits completely in $K$. ∎
2.2. Splitting of primes
We will now review and prove some results on the splitting of primes in non-Galois extensions. The results in this section guarantee that it suffices to assume the existence of one rational prime satisfying the suitable splitting behaviors assumed in Theorem 1.6 and Theorem 1.7, respectively.
Before proving the results, let us first define some necessary notations. For a set $X$, let $\mathrm{Sym}(X)$ denote the symmetric group on $X$. For any $n\in\mathbb{Z}^{+}$, let $S_{n}$ denote the symmetric group of degree $n$, i.e., $S_{n}:=\mathrm{Sym}(\{1,2,\ldots,n\})$. We will also adopt the usual cycle notation, i.e., $(a_{1},a_{2},\ldots,a_{r})$, to represent elements in $S_{n}$.
Lemma 2.3.
Let $K/E$ be a finite extension of number fields and let $\widetilde{K}/E$ be the Galois closure of $K/E$. Let $G=\operatorname{Gal}(\widetilde{K}/E)$, $H=\operatorname{Gal}(\widetilde{K}/K)$, and $\Sigma=\{\sigma:K\rightarrow\widetilde{K}\;\big{|}\;\sigma|_{E}=\operatorname{id}_{E}\}$. Let $G$ (and hence $H$) act on $\Sigma$ via pre-composition. Write $n=[K:E]$. Then,
  1. (1)
    $|\Sigma|=n$ and $G$ acts transitively on $\Sigma$.
  2. (2)
    The action of $G$ on $\Sigma$ induces an embedding $G\hookrightarrow\mathrm{Sym}(\Sigma)$.
  3. (3)
    Let $\sigma_{n}\in\Sigma$ be the unique element such that $\sigma_{n}=\operatorname{id}_{K}$. Then, $H=\operatorname{Stab}_{G}(\{\sigma_{n}\})$ and the action of $H$ on $\Sigma\backslash\{\sigma_{n}\}$ induces an embedding $H\hookrightarrow\mathrm{Sym}(\Sigma\backslash\{\sigma_{n}\})$.
  4. (4)
    $H\backslash G\simeq\Sigma$ as sets with $G$-action.
Remark 2.4.
If we identify $\Sigma$ with $\{1,\ldots,n\}$ such that $\sigma_{n}$ is identified with $n$, then it follows from the lemma that $G$ can be identified with a subgroup of $S_{n}$ and $H=\operatorname{Stab}_{G}(\{n\})$ can be identified with a subgroup of $S_{n}$ consisting of cycles not containing $n$.
Proof.
As $K/E$ is a finite extension of number fields, there exists $\alpha\in K$ such that $K=E(\alpha)$. Let $m_{\alpha}\in E[X]$ be the minimal polynomial of $\alpha$ over $E$. Then, $m_{\alpha}$ is of degree $n$ and has exactly $n$ distinct roots. There is then a one-to-one correspondence between elements $\sigma\in\Sigma$ and roots $\sigma(\alpha)$ of $m_{\alpha}$. Hence, $|\Sigma|=n$. In particular, the action of $G$ on $\Sigma$ corresponds to the action of $G$ on the set of roots of $m_{\alpha}$, which is transitive.
Now, we will show that $G$ acts faithfully on $\Sigma$ so that $G\hookrightarrow\mathrm{Sym}(\Sigma)$. Let $g_{1},g_{2}\in G$ with $g_{1}\sigma=g_{2}\sigma$ for all $\sigma\in\Sigma$. Then, $g_{1}(\sigma(\alpha))=g_{2}(\sigma(\alpha))$ for all $\sigma\in\Sigma$. Hence, $g_{1}=g_{2}$ on all roots of $m_{\alpha}$. Since $\widetilde{K}$ is the splitting field of $m_{\alpha}(X)$ over $E$, it follows that $g_{1}=g_{2}$.
For the third part, it suffices to show that $H=\operatorname{Stab}_{G}(\{\sigma_{n}\})$ so that it follows from the faithful action that $H\hookrightarrow\mathrm{Sym}(\Sigma\backslash\{\sigma_{n}\})$. Note that any element $\sigma\in\Sigma$ is uniquely determined by $\sigma(\alpha)$. Hence, $g\in\operatorname{Stab}_{G}(\{\sigma_{n}\})$ is equivalent to $g(\sigma_{n}(\alpha))=\sigma_{n}(\alpha)$, i.e., $g(\alpha)=\alpha$. As $K=E(\alpha)$, $g(\alpha)=\alpha$ is then equivalent to $g|_{K}=\operatorname{id}_{K}$, which is equivalent to $g\in H$. Hence, $H=\operatorname{Stab}_{G}(\{\sigma_{n}\})$.
The last part simply follows from the fact from group theory that if $G$ acts transitively on $\Sigma$, then $\operatorname{Stab}_{G}(\{\sigma\})\backslash G\simeq\Sigma$ as sets with $G$-action for all $\sigma\in\Sigma$. ∎
Let $K/E$ be an extension of number fields of degree $n$. Given a partition $\mathcal{P}=(a_{1},\ldots,a_{r})$ of $n$, i.e., $n=\sum_{i=1}^{r}a_{i}$ with $a_{i}\in\mathbb{Z}^{+}$ for each $1\leq i\leq r$, we say that a prime $\mathfrak{p}$ of $E$ is of splitting type $\mathcal{P}$ in $K$ if it is unramified in $K$ and $\mathfrak{p}=\wp_{1}\cdots\wp_{r}$ in $K$ where the inertia degree of each $\wp_{i}$ is $a_{i}$. On the other hand, given a partition $\mathcal{P}=(a_{1},\ldots,a_{r})$ of $n$, we say that an element $g\in S_{n}$ is of cycle type $\mathcal{P}$ if it is a product of disjoint cycles of length $a_{1},\ldots,a_{r}$.
Let us also recall two facts from the theory of field extensions. Let $K/E$ and $K^{\prime}/E$ be two finite extensions of number fields. Then, any prime of $E$ that is unramified in both $K$ and $K^{\prime}$ stays unramified in $KK^{\prime}$. In particular, any prime of $E$ that is unramified in $K$ stays unramified in the Galois closure. Also, if $K/E$ and $K^{\prime}/E$ are Galois extensions, then so is $KK^{\prime}/E$.
Lemma 2.5.
Let $L/K/E$ be a tower of finite extensions of number fields such that $L/E$ is a Galois extension. Let $G=\operatorname{Gal}(L/E)$ and $H=\operatorname{Gal}(L/K)$. Let $p$ be a prime of $E$ that is unramified in $L/E$ and let $\mathfrak{p}$ be a prime of $L$ lying above $p$. Let $G_{\mathfrak{p}}=\{\sigma\in G:\sigma(\mathfrak{p})=\mathfrak{p}\}$ and define an action of $G_{\mathfrak{p}}$ on $H\backslash G$ by $H\sigma\cdot g=H\sigma g$. Then,
  1. (1)
    there is a one-to-one correspondence between orbits of $H\backslash G$ under the action of $G_{\mathfrak{p}}$ and primes of $K$ that lie above $p$;
  2. (2)
    the inertia degree of the prime of $K$ corresponding to an orbit of $H\backslash G$ is the size of the orbit.
Proof.
See [Jan96, Proposition 2.7, Chapter III]. ∎
Lemma 2.6.
Let $K/E$ be an extension of number fields of degree $n$ and let $\widetilde{K}/E$ be the Galois closure of $K/E$. Identify $\Sigma=\{\sigma\colon K\rightarrow\widetilde{K}\;:\;\sigma|_{E}=\operatorname{id}_{E}\}$ with $\{1,\ldots,n\}$, $G=\operatorname{Gal}(\widetilde{K}/E)$ with a subgroup of $S_{n}$, and $H=\operatorname{Gal}(\widetilde{K}/K)$ with $\operatorname{Stab}_{G}(\{n\})$. Let $\mathcal{P}=(a_{1},\ldots,a_{r})$ be a partition of $n$. Let $\mathfrak{p}$ be a prime of $E$ that is unramified in $K$ and let $g\in G$ be a Frobenius lift at $\mathfrak{p}$. Then, $g$ is of cycle type $\mathcal{P}$ if and only if $\mathfrak{p}$ is of splitting type $\mathcal{P}$ in $K$.
In particular, $G$ contains an element of cycle type $\mathcal{P}$ if and only if there exists a prime of $E$ that is of splitting type $\mathcal{P}$ in $K$.
Proof.
As $g$ is a Frobenius lift at $\mathfrak{p}$ and $\mathfrak{p}$ is unramified in $K$ and hence in $\widetilde{K}$, we may identify the decomposition group $G_{\mathfrak{p}}\subseteq G$ at $\mathfrak{p}$ with $\langle g\rangle$. Now, $H\backslash G\simeq\Sigma\simeq\{1,\ldots,n\}$ as sets with $G$-action by Lemma 2.3. In order to obtain the splitting behavior of $\mathfrak{p}$ in $K/E$, by Lemma 2.5, it suffices to look at the sizes of orbits of $\{1,\ldots,n\}$ under the action of $G_{\mathfrak{p}}=\langle g\rangle$. It is now easy to see that the following are equivalent:
  1. (i)
    $g$ is of cycle type $\mathcal{P}$;
  2. (ii)
    the sizes of orbits of $\{1,\ldots,n\}$ under the action of $G_{\mathfrak{p}}=\langle g\rangle$ are exactly given by $\mathcal{P}$;
  3. (iii)
    $\mathfrak{p}$ is of splitting type $\mathcal{P}$.
Proposition 2.7.
Let $K_{1}/E,\ldots,K_{m}/E$ be finite extensions of number fields with degree $n_{1},\ldots,n_{m}$ respectively. Let $\mathcal{P}_{i}$ be a partition of $n_{i}$. Then, the following are equivalent:
  1. (i)
    there exists a prime of $E$ that is of splitting type $\mathcal{P}_{i}$ in $K_{i}$ for all $i$;
  2. (ii)
    the set of primes of $E$ that are of splitting type $\mathcal{P}_{i}$ in $K_{i}$ for all $i$ has positive density.
Proof.
Clearly, (ii) implies (i), so suppose that (i) holds. Let $\widetilde{K_{i}}/E$ be the Galois closure of $K_{i}/E$ and let $L=\widetilde{K_{1}}\cdots\widetilde{K_{m}}$ be the composite field. Note that $L/E$ is Galois since each $\widetilde{K_{i}}/E$ is Galois. Let $G_{i}=\operatorname{Gal}(\widetilde{K_{i}}/E)\hookrightarrow S_{n_{i}}$ and let $G=\operatorname{Gal}(L/E)$. Let $\pi_{i}:G\twoheadrightarrow G_{i}$ denote the projection map. Note that
$\Sigma_{i}:=\{\sigma\colon K_{i}\rightarrow L\;:\;\sigma|_{E}=\operatorname{id}_{E}\}=\{\sigma\colon K_{i}\rightarrow\widetilde{K_{i}}\;:\;\sigma|_{E}=\operatorname{id}_{E}\},$
and the map $G\rightarrow S_{n_{i}}$ given by the action of $G^{\prime}_{i}$ on $\Sigma_{i}$ factors as $G\twoheadrightarrow G_{i}\hookrightarrow S_{n_{i}}$.
Let $\mathfrak{p}$ be a prime of $E$ that is of splitting type $\mathcal{P}_{i}$ in $K_{i}$ for each $i$. Since $\mathfrak{p}$ is unramified in $K_{i}$ for all $i$, $\mathfrak{p}$ is also unramified in $L$. Let $G_{\mathfrak{p}}=\langle g\rangle\subseteq G$ be the decomposition group at $\mathfrak{p}$, where $g$ is a Frobenius lift at $\mathfrak{p}$, and let $C_{g}$ be the conjugacy class of $g$ in $G$. By the Chebotarev density theorem, there exists a positive density set of primes of $E$ that are unramified in $L/E$ and whose Frobenius elements lie in $C_{g}$. Let $\mathfrak{q}$ be one of them and we may identify $G_{\mathfrak{q}}$ with $\langle g\rangle$ and the Frobenius lift at $\mathfrak{q}$ with $g$. By Lemma 2.6, since $\mathfrak{p}$ is of splitting type $\mathcal{P}_{i}$ in $K_{i}$, each $\pi_{i}(g)\in G_{i}$ is of cycle type $\mathcal{P}_{i}$. It then follows from Lemma 2.6 again that $\mathfrak{q}$ is of splitting type $\mathcal{P}_{i}$ in $K_{i}$, whence (ii) follows. ∎
2.3. Extensions of degree $q$ or $2q$
We will now focus on the splitting behaviors of primes in field extensions of degree $q$ or $2q$ for some prime $q$, which is one of the ingredients for the proof of Theorem 1.1 and Theorem 1.5. We will first prove a lemma on transitive subgroups of $S_{2q}$ and then prove the important proposition on field extensions of degree $q$ or $2q$.
Lemma 2.8.
Let $q$ be a prime and $G\hookrightarrow S_{2q}$ be a transitive subgroup, i.e., $G$ acts transitively on $\{1,2,\ldots,2q\}$. Then, $G$ contains an element that is a product of two disjoint cycles of length $q$.
Proof.
Let $G\cdot\{2q\}$ be the $G$-orbit of $2q$. Note that $|G\cdot\{2q\}|=2q$. Since $G$ is transitive, $2q$ divides $|G|$ by the orbit-stabilizer theorem. Also, $|G|$ divides $|S_{2q}|=(2q)!$. Hence, $1\leq v_{q}(|G|)\leq 3$.
Case 1: $v_{q}(|G|)=3$. Then, $q=2$ and $G$ is a transitive subgroup of $S_{4}$ with $8$ dividing $|G|$. By the classification of subgroups of $S_{4}$, $G$ is either $S_{4}$ or conjugate to $\langle(1,2,3,4),(1,3)\rangle$ in $S_{4}$. In both cases, the element $(1,3)(2,4)$ (or its conjugate in $S_{4}$) lies in $G$, which satisfies the requirement.
Case 2: $v_{q}(|G|)=2$. By Sylow theorems, there exists a Sylow $q$-subgroup $I$ in $G$ of order $q^{2}$. If $q=2$, then $I$ can only be conjugate to $\langle(1,2,3,4)\rangle$, $\langle(1,2),(3,4)\rangle$, or $\langle(1,2)(3,4),(1,3)(2,4)\rangle$ in $S_{4}$. In particular, $I$ always contains an element which is a product of two disjoint cycles of length $q$. Now, suppose that $q\neq 2$. Note that $\langle(1,2,\ldots,q),(q+1,\ldots,2q)\rangle$ is a Sylow $q$-subgroup in $S_{2q}$ of order $q^{2}$, so $I$ is conjugate to $\langle(1,2,\ldots,q),(q+1,\ldots,2q)\rangle$ in $S_{2q}$ by Sylow theorems. Hence, $I$ is generated by two single cycles of length $q$ with disjoint elements. In particular, $I$ contains an element which is a product of two disjoint cycles of length $q$.
Case 3: $v_{q}(|G|)=1$. As $q$ is a prime, $G$ contains an element $\tau$ of order $q$. Note that an element of order $q$ in $S_{2q}$ is either a single cycle of length $q$ or a product of two disjoint cycles of length $q$. Suppose for the sake of contradiction that $\tau$ is a single cycle of length $q$. Then, there exists $a\in\{1,2,\cdots,2q\}$ such that $\tau(a)=a$. Since $G$ is transitive, there exists $\eta\in G$ such that $\eta(a)=2q$. Then, $(\eta\tau\eta^{-1})(2q)=2q$, so $\eta\tau\eta^{-1}\in\operatorname{Stab}_{G}(\{2q\})=:H$. By the orbit-stabilizer theorem, we have $|G\cdot\{2q\}|\cdot|H|=|G|$, so $q\nmid|H|$ as $v_{q}(|G|)=1$. This contradicts that $H$ contains the element $\eta\tau\eta^{-1}$ of order $q$. ∎
Proposition 2.9.
Let $q$ be a prime and let $K/E$ be a finite extension of number fields. Then,
  1. (1)
    if $[K:E]=q$, then there exists a prime of $E$ that is inert in $K$;
  2. (2)
    if $[K:E]=2q$, then there exists a prime of $E$ that splits into two distinct primes in $K$ of the same inertia degree.
Proof.
[…]
Corollary 2.10.
Let $F/\mathbb{Q}$ be a quadratic extension and $K/\mathbb{Q}$ be an extension of degree $q$ for some rational prime $q\geq 3$. Then, there exists a rational prime that splits completely in $F$ and remains inert in $K$.
Proof.
Since $q$ is an odd prime, the composite field $FK$ is of degree $2q$. By Proposition 2.9, there exists a rational prime $p$ that splits into two distinct primes in $FK$ of the same inertia degree $q$. Then, $p$ splits into at most two primes in $K$. Suppose that $p=\mathfrak{p}_{1}\mathfrak{p}_{2}$ in $K$ of inertia degree $f_{1}$ and $f_{2}$, respectively. By the splitting behavior of $p$ in $FK$, both $\mathfrak{p}_{1}$ and $\mathfrak{p}_{2}$ remain inert in $FK$, so their inertia degrees as primes in $FK$ are $2f_{1}$ and $2f_{2}$, respectively. This contradicts that the inertia degree of the two primes in $FK$ that lie above $p$ is an odd prime $q$. Hence, $p$ must remain inert in $K$, and as $p$ splits into two primes in $FK$, $p$ must split completely in $F$ since $F/\mathbb{Q}$ is quadratic. The result then follows. ∎
3. Results on Galois representations
3.1. Galois representations of abelian varieties
In this subsection, we will introduce the general theory of $\ell$-adic and $\lambda$-adic Galois representations of abelian varieties with extra endomorphisms. We will then specialize to the case of “$\operatorname{GL}_{2}$-type” abelian varieties, prove a result on the images of their Galois representations, and give a criterion, distinct from the definition, to identify their ordinary primes.
[…]
$\rho_{A,\ell}:\operatorname{Gal}(\overline{F}/F)\to\operatorname{Aut}(V_{\ell}(A))\simeq\operatorname{GL}_{2g}(\mathbb{Q}_{\ell}),$
which we call the $\ell$-adic Galois representation of $A$ [^1] Due to the symplectic pairing on $V_{\ell}(A)$ coming from the polarization of $A$, the image of $\rho_{A,\ell}$ actually lies in $\operatorname{GSp}_{2g}(\mathbb{Q}_{\ell})$. This Galois representation satisfies that for all primes $v$ of $F$ with $v\nmid N_{A}\ell$, $\rho_{A,\ell}$ is unramified at $v$, and the characteristic polynomial
$P_{A,v}(X):=\operatorname{det}(XI-\rho_{A,\ell}(\operatorname{Frob}_{v}))$ (1)
is a degree $2g$ polynomial with integer coefficients, which is independent of the choice of $\ell$. By the Weil conjectures, for each embedding $\iota:K\rightarrow\mathbb{C}$ and each root $\alpha$ of $P_{A,v}(X)$, we have $|\iota(\alpha)|_{\mathbb{C}}=\mathrm{N}(v)^{1/2}$.
Now, we will discuss the case when $A$ has extra endomorphisms. The main reference for this part is [Rib76, II]. Let $\operatorname{End}_{F}(A)$ denote the ring of $F$-endomorphisms of $A$ and let $\operatorname{End}_{\overline{F}}(A)$ denote the ring of geometric endomorphisms of $A$. Assume that $\operatorname{End}_{F}(A)\otimes\mathbb{Q}$ contains a number field $K$. Then, it is known that $n=\frac{2g}{[K:\mathbb{Q}]}$ is necessarily an integer. In this case, we say that $A/F$ is of $\operatorname{GL}_{n}(K)$-type or simply of $\operatorname{GL}_{n}$-type without referring to the number field.
Suppose from now on that $A/F$ is of $\operatorname{GL}_{n}(K)$-type. Then, the extra endomorphisms induce more restrictions on the Galois image. Specifically, if we set $K_{\ell}:=K\otimes\mathbb{Q}_{\ell}$, then the Galois action on $V_{\ell}(A)$ is $K_{\ell}$-linear, and moreover $V_{\ell}(A)$ is a free $K_{\ell}$-module of rank $n$. Hence, the $\ell$-adic Galois representation can be interpreted as
$\rho_{A,\ell}:\operatorname{Gal}(\overline{F}/F)\to\operatorname{Aut}_{K_{\ell}}(V_{\ell}(A))\simeq\operatorname{GL}_{n}(K_{\ell})=\operatorname{GL}_{n}(K\otimes\mathbb{Q}_{\ell}).$
Since $K_{\ell}=\prod_{\lambda|\ell}K_{\lambda}$, for each prime $\lambda|\ell$ of $K$, we can put
$V_{\lambda}(A):=V_{\ell}(A)\otimes_{K_{\ell}}K_{\lambda}.$
Then, for each prime $\lambda$ of $K$, $V_{\lambda}(A)$ is an $n$-dimensional $K_{\lambda}$-vector space with the induced action of $\operatorname{Gal}(\overline{F}/F)$. Hence, each $V_{\lambda}(A)$ induces a continuous Galois representation
$\rho_{A,\lambda}:\operatorname{Gal}(\overline{F}/F)\to\operatorname{Aut}_{K_{\lambda}}(V_{\lambda}(A))\simeq\operatorname{GL}_{n}(K_{\lambda}),$
which we call the $\lambda$-adic Galois representation of $A$. Alternatively, $\rho_{A,\lambda}$ can be viewed as the composition of $\rho_{A,\ell}$ and the projection map $\operatorname{GL}_{n}(K\otimes\mathbb{Q}_{\ell})\rightarrow\operatorname{GL}_{n}(K_{\lambda})$. The Galois representations $\rho_{A,\ell}$ (under the new interpretation) and $\rho_{A,\lambda}$ (with $\lambda|\ell$) satisfy that for all primes $v$ of $F$ with $v\nmid N_{A}\ell$, $\rho_{A,\ell}$ and $\rho_{A,\lambda}$ are unramified at $v$, and the characteristic polynomials
$\operatorname{det}(XI-\rho_{A,\ell}(\operatorname{Frob}_{v}))\in(\mathcal{O}_{K}\otimes\mathbb{Z})[X]\quad\text{ and }\quad\operatorname{det}(XI-\rho_{A,\lambda}(\operatorname{Frob}_{v}))\in\mathcal{O}_{K}[X]$
are of degree $n$ and independent of $\ell$ and $\lambda$, where we used the interpretation $\rho_{A,\ell}:\operatorname{Gal}(\overline{F}/F)\to\operatorname{GL}_{n}(K\otimes\mathbb{Q}_{\ell})$. Moreover, it is known by [Shi67, Proposition 11.9, p. 586] (see also [Chi92, Proposition 2.2, p. 322]) that
$P_{A,v}(X)=\prod_{\sigma:K\rightarrow\mathbb{C}}\sigma\big{(}\operatorname{det}(XI-\rho_{A,\lambda}(\operatorname{Frob}_{v}))\big{)}.$ (2)
Let us now specialize to the case when $n=2$, i.e., when $A/F$ is of $\operatorname{GL}_{2}(K)$-type. For a prime $v$ of $F$ with $v\nmid N_{A}$ (and $v\nmid\mathrm{N}(\lambda)$), we write
$a_{v}(A)=\operatorname{tr}\rho_{{A,\lambda}}(\operatorname{Frob}_{v})\quad\text{ and }\quad d_{v}(A)=\operatorname{det}\rho_{{A,\lambda}}(\operatorname{Frob}_{v})$
so that
$\operatorname{det}(XI-\rho_{A,\lambda}(\operatorname{Frob}_{v}))=X^{2}-a_{v}(A)X+d_{v}(A)\in\mathcal{O}_{K}[X].$
To summarize, we have a family of Galois representations $\{\rho_{A,\ell}:\operatorname{Gal}(\overline{F}/F)\rightarrow\operatorname{GL}_{2}(K\otimes\mathbb{Q}_{\ell})\}_{\ell\in\Sigma_{\mathbb{Q}}}$ such that for all rational primes $\ell$ and all primes $v$ of $F$ with $v\nmid N_{A}\ell$, $\rho_{A,\ell}$ is unramified at $v$, the characteristic polynomial of the Frobenius satisfies that
$\operatorname{det}(XI-\rho_{A,\ell}(\operatorname{Frob}_{v}))=X^{2}-(a_{v}(A)\otimes 1)X+(d_{v}(A)\otimes 1)\in(\mathcal{O}_{K}\otimes\mathbb{Z})[X]$
is of degree $2$ and independent of $\ell$, and for all embeddings $\iota:K\rightarrow\mathbb{C}$,
$|\iota(a_{v}(A))|_{\mathbb{C}}\leq 2\sqrt{\mathrm{N}(v)}.$
The last pieces of information about the Galois representations of $A$ needed for the proof of Theorem 1.3 are the determinant of $\rho_{A,\lambda}$ and image of $\rho_{A,\ell}$, which we present now.
The analysis of the shapes of the determinant and the Galois image depends heavily on [FFG24, Proposition 6.10] and [BGK06, Theorem 6.16], and we will recall the former one here as we will need it in several places. We will also recall several definitions from [FFG24] and [BGK06] to make this part self-contained. Indeed, similar result with extra restrictions on the endomorphism algebra of $A$ was already established by Ribet [Rib76, Theorem 5.5.2, p. 801] and reformulated by Lombardo [Lom16, Theorem 1.4 and Remark 1.6].
Definition 3.1 ([FFG24]).
We say that an abelian variety $A/F$ is of $\operatorname{GL}_{n}$-type if there exists a number field $K\hookrightarrow\operatorname{End}_{F}(A)\otimes\mathbb{Q}$ with $n=\frac{2\dim A}{[K:Q]}$. Moreover, we say that $A/F$ is genuinely of $\operatorname{GL}_{n}$-type if it is simple, of $\operatorname{GL}_{n}$-type, and its base change to $\overline{F}$ has no simple subvariety of $\operatorname{GL}_{m}$-type with $m<n$.
Definition 3.2 ([FFG24]).
We say that a simple abelian variety $A/F$ is of the first kind if $A$ is of type I, II, or III in Albert’s classification,  [^3] For a detailed summary of the Albert’s classification, see [Mum70, pp. 201–202]. or equivalently, the center of $\operatorname{End}_{F}(A)\otimes\mathbb{Q}$ is a totally real field. For an abelian variety $A/F$ that is genuinely of $\operatorname{GL}_{n}$-type, its base change to $\overline{F}$ is necessarily isogenous to $B^{r}$ for some simple abelian variety $B/\overline{F}$, and we say that $A/F$ is geometrically of the first kind if $B$ is of the first kind.
Definition 3.3 ([BGK06]).
We say that an abelian variety $A/F$ of dimension $g$ is of class $\mathcal{A}$ if
  1. (1)
    $A/F$ is a simple, principally polarized abelian variety.
  2. (2)
    $A/\overline{F}$ is of type I or II in Albert’s classification and $\operatorname{End}_{F}(A)\otimes\mathbb{Q}=\operatorname{End}_{\overline{F}}(A)\otimes\mathbb{Q}$.
  3. (3)
    For every rational prime $\ell$, the Zariski closure of the $\rho_{A,\ell}(\operatorname{Gal}(\overline{F}/F))$ in $\operatorname{GL}_{2g}/\mathbb{Q}_{\ell}$ is a connected algebraic group.
  4. (4)
    $g=ehd$ where $h$ is an odd integer, $e$ is the degree of the center $E$ of $D$, and $d^{2}=[D:E]$.
Proposition 3.4 ([FFG24, Proposition 6.10]).
Let $A/F$ be an abelian variety that is genuinely of $\operatorname{GL}_{2}$-type and geometrically of the first kind. Then, there exists a finite order character $\epsilon:\operatorname{Gal}(\overline{F}/F)\rightarrow K^{\times}$ such that for all primes $\lambda$ of $K$,
$\operatorname{det}\rho_{A,\lambda}=\epsilon\chi_{\ell},$
where $\ell$ is the rational prime lying below $v$ and $\chi_{\ell}:\operatorname{Gal}(\overline{F}/F)\rightarrow K_{\lambda}^{\times}$ denotes the $\ell$-adic cyclotomic character. Moreover, if the center of $\operatorname{End}_{F}(A)\otimes\mathbb{Q}$ is a totally real field, then $\operatorname{det}\rho_{A,\lambda}=\chi_{\ell}$.
With all the tools in hand, we are now ready to prove the results on the determinant and the Galois image.
Lemma 3.5.
Let $A/F$ be an abelian variety of dimension $g$ and conductor $N_{A}$ over a number field $F$. Assume that $A$ is absolutely simple, $\operatorname{End}_{F}(A)\otimes\mathbb{Q}$ contains a number field $K$ of degree $g$, and $A/\overline{F}$ is of the first kind. Then, there exists a finite order character $\epsilon:\operatorname{Gal}(\overline{F}/F)\rightarrow K^{\times}$ such that for all rational primes $\ell$, all primes $\lambda$ of $K$ with $\lambda|\ell$, and all primes $v$ of $F$ with $v\nmid N_{A}\ell$,
$\operatorname{det}(XI-\rho_{A,\lambda}(\operatorname{Frob}_{v}))$ $=X^{2}-a_{v}(A)X+\epsilon(v)\mathrm{N}(v)\in\mathcal{O}_{K}[X]$
$\operatorname{det}(XI-\rho_{A,\ell}(\operatorname{Frob}_{v}))$ $=X^{2}-(a_{v}(A)\otimes 1)X+(\epsilon(v)\mathrm{N}(v)\otimes 1)\in(\mathcal{O}_{K}\otimes\mathbb{Z})[X],$
where $\rho_{A,\ell}$ is interpreted as $\operatorname{Gal}(\overline{F}/F)\rightarrow\operatorname{GL}_{2}(K\otimes\mathbb{Q}_{\ell})$.
Proof.
It is easy to check that the assumptions on $A$ guarantee that $A$ is genuinely of $\operatorname{GL}_{2}$-type and geometrically of the first kind. The result then simply follows from Proposition 3.4 (and our previous discussion). ∎
Theorem 3.6.
Let $A/F$ be an abelian variety of dimension $g$ over a number field $F$. Assume that $A$ is absolutely simple, $\operatorname{End}_{\overline{F}}(A)\otimes\mathbb{Q}$ contains a number field $K$ of degree $g$, [^4] Note that this is weaker than the assumption $K\hookrightarrow\operatorname{End}_{F}(A)\otimes\mathbb{Q}$ in Theorem 1.3. and the center of $\operatorname{End}_{\overline{F}}(A)\otimes\mathbb{Q}$ is a totally real field $K^{\prime}$. Then, there exists a finite extension $F^{\prime}/F$ such that for all sufficiently large prime $\ell$, we have, up to conjugation, that
$\rho_{A,\ell}(\operatorname{Gal}(\overline{F^{\prime}}/F^{\prime}))=\{M\in\operatorname{GL}_{2}(\mathcal{O}_{K^{\prime}}\otimes\mathbb{Z}_{\ell}):\operatorname{det}M\in\mathbb{Z}_{\ell}^{\times}\}.$
Proof.
Let $F^{\prime}/F$ be a finite extension of number fields such that  [^5] The existence of this finite extension can be deduced from [Sil92, Theorem 4.1] and [SZ95, Lemma 2.7 and Remark 3.1] (see also [Ser00]).
  1. (1)
    $A$ is principally polarized over $F^{\prime}$;
  2. (2)
    $D:=\operatorname{End}_{F^{\prime}}(A)\otimes\mathbb{Q}=\operatorname{End}_{\overline{F}}(A)\otimes\mathbb{Q}$;
  3. (3)
    for every rational prime $\ell$, the Zariski closure of $\rho_{A,\ell}(\operatorname{Gal}(\overline{F}/F^{\prime}))$ in the algebraic group $\operatorname{GL}_{2g}/\mathbb{Q}_{\ell}$ is a connected algebraic group.
By the assumptions and (2), $A/F^{\prime}$ is simple and of $\operatorname{GL}_{2}$-type, and by [Wu20, Proposition 1.7], $A/F^{\prime}$ is not of type III. As the center $K^{\prime}$ of $D$ is assumed to be totally real, by [Wu20, Proposition 1.7], we have either $D=K^{\prime}(=K)$ or $D$ is a division quaternion algebra over $K^{\prime}$ with $[K^{\prime}:\mathbb{Q}]=g/2$. Putting all things together, it is then easy to check that $A/F^{\prime}$ is of class $\mathcal{A}$, so by [BGK06, Theorem 6.16 (6.18)], we have $\operatorname{SL}_{2}(\mathcal{O}_{K^{\prime}}\otimes\mathbb{Z}_{\ell})\subseteq\rho_{A,\ell}(\operatorname{Gal}(\overline{F^{\prime}}/F^{\prime}))$ for sufficiently large prime $\ell$. On the other hand, by the assumptions, $A/F^{\prime}$ is (geometrically) of the first kind. As the center $K^{\prime}$ of $D$ is totally real, by Proposition 3.4 (and gathering all primes $\lambda^{\prime}$ of $K^{\prime}$ with $\lambda^{\prime}|\ell$), we have $\operatorname{det}\rho_{A,\ell}|_{\operatorname{Gal}(\overline{F}/F^{\prime})}=\chi_{\ell}$ for all rational primes $\ell$.
Finally, for sufficiently large prime $\ell$, since
$\rho_{A,\ell}(\operatorname{Gal}(\overline{F^{\prime}}/F^{\prime}))\subseteq\{M\in\operatorname{GL}_{2}(\mathcal{O}_{K^{\prime}}\otimes\mathbb{Z}_{\ell}):\operatorname{det}M\in\mathbb{Z}_{\ell}^{\times}\}$ (3)
and the exact sequence
$0\to\operatorname{SL}_{2}(\mathcal{O}_{K^{\prime}}\otimes\mathbb{Z}_{\ell})\to\{M\in\operatorname{GL}_{2}(\mathcal{O}_{K^{\prime}}\otimes\mathbb{Z}_{\ell}):\operatorname{det}M\in\mathbb{Z}_{\ell}^{\times}\}\xrightarrow{\operatorname{det}}\mathbb{Z}_{\ell}^{\times}\to 0$
also holds for $\rho_{A,\ell}(\operatorname{Gal}(\overline{F^{\prime}}/F^{\prime}))$, the inclusion (3) is actually an equality by the five lemma (see also [Rib75, Corollary 2.2]). ∎
To end this subsection, we will give a different criterion of ordinary primes of $A$ using the Frobenius trace $a_{v}(A)$ of the $\lambda$-adic representations. Here, we take one of the equivalent definitions of ordinary primes from [Fit24].
Definition 3.7.
Let $A$ be an abelian variety of dimension $g$ and conduct $N_{A}$ defined over a number field $F$. We say that a prime $v$ of $F$ is an ordinary prime of $A/F$ if $v\nmid N_{A}$ and $p\nmid a_{v,g}(A)$, where $p$ is the rational prime lying below $v$ and $a_{v,g}(A)$ is the coefficient of the term $X^{g}$ of $P_{A,v}(X)$ defined by Equation 1.
Proposition 3.8.
Let $A/F$ be an abelian variety of dimension $g$ and conductor $N_{A}$ over a number field $F$. Assume that $A$ is absolutely simple, $\operatorname{End}_{F}(A)\otimes\mathbb{Q}$ contains a number field $K$ of degree $g$, and $A/\overline{F}$ is of the first kind. Then, a prime $v\nmid N_{A}$ of $F$ is an ordinary prime of $A$ if and only if $\lambda\nmid(a_{v}(A))$ for all primes $\lambda|\mathrm{N}(v)$ of $K$.
Proof.
It is an easy fact from algebraic number theory that for a rational prime $p$, $p\nmid\mathrm{N}_{K/\mathbb{Q}}(a_{v}(A))$ if and only if $\lambda\nmid(a_{v}(A))$ for all primes $\lambda|p$ of $K$. We will thus prove that $v$ is an ordinary prime of $A$ if and only if $p\nmid\mathrm{N}_{K/\mathbb{Q}}(a_{v}(A))$ for the rational prime $p$ lying below $v$.
By Equation 2,
$P_{A,v}(X)=\prod_{\sigma:K\rightarrow\mathbb{C}}(X^{2}-\sigma(a_{v}(A))X+\sigma(\epsilon(v))\mathrm{N}(v)).$
Since #$\#\{\sigma:K\rightarrow\mathbb{C}\}=g$, it is easy to see that the coefficient $a_{v,g}(A)$ of the term $X^{g}$ is of the form
$a_{v,g}(A)=\prod_{\sigma:K\rightarrow\mathbb{C}}\sigma(a_{v}(A))+\mathrm{N}(v)\alpha=\mathrm{N}_{K/\mathbb{Q}}(a_{v}(A))+\mathrm{N}(v)\alpha$
for some algebraic integer $\alpha$. Note that both $a_{v,g}(A)$ and $\mathrm{N}_{K/\mathbb{Q}}(a_{v}(A))$ are integers, so $\alpha\in\mathbb{Q}$, and hence $\alpha\in\mathbb{Z}$. In particular, $p\nmid a_{v,g}(A)$ if and only if $p\nmid\mathrm{N}_{K/\mathbb{Q}}(a_{v}(A))$. The result then follows. ∎
3.2. Galois representations of modular forms
In this subsection, we will introduce the $\ell$-adic Galois representations of modular forms and recall the properties of the images of those Galois representations. We will also briefly review the Eichler–Shimura theory and relate the $p$-ordinariness of weight $2$ modular forms to the $p$-ordinariness of the associated abelian varieties.
Let $f\in S_{k}(\Gamma_{0}(N),\epsilon)$ be a newform of weight $k$, level $N$, and nebentypus $\epsilon$, i.e., $f$ is a normalized Hecke eigenform which is new in level $N$ and $\epsilon:\mathbb{Z}\rightarrow\mathbb{C}$ is any Dirichlet character of conductor dividing $N$. Let $a_{n}(f)$ be the $n$-th Fourier coefficient of $f$ and let $K_{f}=\mathbb{Q}(\{a_{n}(f):(n,N)=1\})$ be the coefficient field of $f$. It is known that each $a_{n}(f)$ is an algebraic integer, $K_{f}$ is a number field, and $\mathcal{O}_{K_{f}}$ contains the image of $\epsilon$. Let $\ell$ be any rational prime. By the works of Eichler–Shimura for $k=2$, Deligne for $k\geq 2$ [Del71], and Deligne–Serre for $k=1$ [DS74], there exists a continuous Galois representation
$\rho_{f,\ell}:\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})\rightarrow\operatorname{GL}_{2}(K_{f}\otimes\mathbb{Q}_{\ell})$
such that for every rational primes $p\nmid\ell N$, $\rho_{f,\ell}$ is unramified at $p$ and
$\operatorname{det}(XI-\rho_{f,\ell}(\operatorname{Frob}_{p}))=X^{2}-(a_{p}(f)\otimes 1)X+(\epsilon(p)p^{k-1}\otimes 1)\in(\mathcal{O}_{K_{f}}\otimes\mathbb{Z})[X].$
By Ramanujan’s conjecture, it is known that $|\iota(a_{p}(f))|_{\mathbb{C}}\leq 2p^{\frac{k-1}{2}}$ for all embeddings $\iota:K\hookrightarrow\mathbb{C}$.
As in the case of abelian varieties, the image of $\rho_{f,\ell}$ is impacted by the “inner twist” of $f$, which we introduce now. The main references for this part are [Rib80] and [Rib85]. The modular form $f$ is said to have complex multiplication (CM) or called a CM modular form if there exists a nontrivial Dirichlet character $\chi$ such that $a_{p}(f)=\chi(p)a_{p}(f)$ for all but finitely many primes $p$. More generally, the modular form $f$ is said to have inner twist if there exists a nontrivial Dirichlet character $\chi$ and an embedding $\sigma:K_{f}\rightarrow\mathbb{C}$ such that
$\sigma(a_{p}(f))=\chi(p)a_{p}(f)$
for all but finitely many primes $p$. We also say that $f$ is a non-CM modular form if $f$ has no complex multiplication.
Suppose that $f$ is a non-CM modular form. Then, for each embedding $\sigma:K_{f}\rightarrow\mathbb{C}$, there exists at most one Dirichlet character $\chi=\chi_{\sigma}$ such that $f$ has inner twist by $(\sigma,\chi)$. All such embeddings $\sigma$ form a group $\Gamma_{f}$ and we set
$H_{f}:=\bigcap_{\sigma\in\Gamma_{f}}\mathrm{Ker}\chi_{\sigma},\quad F_{f}^{\prime}:=\overline{\mathbb{Q}}^{H_{f}},\quad K_{f}^{\prime}:=K_{f}^{\Gamma_{f}}.$
Consequently, we have the following big image theorem for $\rho_{f,\ell}$, where $f$ possibly has inner twist.
Theorem 3.9 ([Rib85, Theorem 3.1]).
We keep the notation above. Let $k\geq 2$ and $f\in S_{k}(\Gamma_{0}(N),\epsilon)$ be a non-CM newform. Then for all sufficiently large prime $\ell$, we have, up to conjugation, that
$\rho_{f,\ell}(\operatorname{Gal}(\overline{\mathbb{Q}}/F_{f}^{\prime}))=\big{\{}M\in\operatorname{GL}_{2}(\mathcal{O}_{K_{f}^{\prime}}\otimes\mathbb{Z}_{\ell}):\operatorname{det}M\in(\mathbb{Z}_{\ell}^{\times})^{k-1}\big{\}}.$
We will now relate the ordinary primes of modular forms (of weight $2$) to the ones of abelian varieties. Let us first recall our definition of ordinary primes for modular forms.
Definition 3.10.
Let $f$ be a newform of weight $k\geq 2$ and level $N$, let $a_{n}(f)$ be the $n$-th Fourier coefficient of $f$, and let $K_{f}$ be the coefficient field of $f$. For a rational prime $p\nmid N$ and a prime $\lambda|p$ of $K_{f}$, $f$ is called $\lambda$-ordinary if $\lambda\nmid(a_{p}(f))$, and $f$ is called $p$-ordinary if $f$ is $\lambda$-ordinary for all $\lambda|p$, or equivalently, $\lambda\nmid(a_{p}(f))$ for all primes $\lambda|p$ of $K_{f}$.
Remark 3.11.
It is an easy fact from algebraic number theory that for any number field $K$, $\alpha\in\mathcal{O}_{K}$, and rational prime $p$, $p|\mathrm{N}_{K/\mathbb{Q}}(\alpha)$ if and only if $\lambda|(\alpha)$ for some prime $\lambda|p$ of $K$. From this, one can give another equivalent definition that for a rational prime $p\nmid N$, $f$ is called $p$-ordinary if $p\nmid\mathrm{N}_{K_{f}/\mathbb{Q}}(a_{p}(f))$.
Let us now specialize to the case when $k=2$, i.e., $f\in S_{2}(\Gamma_{0}(N),\epsilon)$ is a newform of weight $2$. By the Eichler–Shimura theory, one can associate $f$ with an abelian variety $A_{f}/\mathbb{Q}$  [^6] This abelian variety $A_{f}/\mathbb{Q}$ is constructed by taking a quotient of the Jacobian of the modular curve $X_{1}(N)$, and we refer to [DS05, Section 6.6] and [Roh97] for a detailed review of the Eichler–Shimura theory. of dimension $g=[K_{f}:\mathbb{Q}]$ such that $K_{f}=\operatorname{End}_{\mathbb{Q}}(A_{f})\otimes\mathbb{Q}$ and that for all rational primes $p\nmid N$, $A_{f}$ has good reduction at $p$ and
$P_{A_{f},p}(X)=\prod_{\sigma:K_{f}\rightarrow\mathbb{C}}\big{(}X^{2}-\sigma(a_{p}(f))X+\sigma(\epsilon(p))p\big{)},$
where $P_{A_{f},p}(X)$ is the polynomial defined in Equation 1.
Proposition 3.12.
Let $f\in S_{2}(\Gamma_{0}(N),\epsilon)$ be a newform and let $A_{f}/\mathbb{Q}$ be the associated abelian variety. Then, for any rational prime $p\nmid N$, $f$ is ordinary at $p$ if and only if $A_{f}$ is ordinary at $p$.
Proof.
Note that $f$ is ordinary at $p$ if and only if $p\nmid\mathrm{N}_{K_{f}/\mathbb{Q}}(a_{p}(f))$. We will prove that the latter is equivalent to $A_{f}$ being ordinary at $p$. For all primes $p\nmid N$,
$P_{A_{f},p}(X)=\prod_{\sigma:K_{f}\rightarrow\mathbb{C}}\big{(}X^{2}-\sigma(a_{p}(f))X+\sigma(\epsilon(p))p\big{)}.$
Since #$\#\{\sigma:K_{f}\rightarrow\mathbb{C}\}=[K_{f}:\mathbb{Q}]=:g$, it is easy to see that the coefficient $a_{p,g}(A_{f})$ of the term $X^{g}$ is of the form
$a_{p,g}(A_{f})=\prod_{\sigma:K_{f}\rightarrow\mathbb{C}}\sigma(a_{p}(f))+p\alpha=\mathrm{N}_{K_{f}/\mathbb{Q}}(a_{p}(f))+p\alpha$
for some algebraic integer $\alpha$. Note that both $a_{p,g}(A_{f})$ and $\mathrm{N}_{K_{f}/\mathbb{Q}}(a_{p}(f))$ are integers, so $\alpha\in\mathbb{Q}$, and hence $\alpha\in\mathbb{Z}$. In particular, $p\nmid a_{p,g}(A_{f})$ if and only if $p\nmid\mathrm{N}_{K_{f}/\mathbb{Q}}(a_{p}(f))$. The result then follows. ∎
4. Proof of Theorem 1.7
We will first give a proof of Theorem 1.7 since its proof is much easier. We will adopt all the notations in Theorem 1.7 unless stated otherwise. Recall that in this case, we have a rational prime $p$ that lies below a degree $1$ prime of $F$ and remains inert in $K$.
Lemma 4.1.
Let $F$ and $K$ be two number fields and let $S$ be a finite set of primes of $F$. Let $(a_{v})_{v\in\Sigma_{F}\setminus S}$ be a fixed sequence in $\mathcal{O}_{K}$ satisfying that
$|\iota(a_{v})|_{\mathbb{C}}\leq 2\mathrm{N}(v)^{1/2}$
for all $v\in\Sigma_{F}$ with $v\notin S$ and all embeddings $\iota:K\rightarrow\mathbb{C}$. Let $v\notin S$ be a degree $1$ prime of $F$ and let $p$ be the rational prime lying below $v$. Suppose that $p\geq 5$ and $p$ is inert in $K$. Then, $a_{v}=0$ or $v\in S^{\mathrm{ord}}$.
Proof.
Suppose that $v\notin S^{\mathrm{ord}}$. Then, there exists a prime $\lambda$ of $K$ with $\lambda|\mathrm{N}(v)$ such that $\lambda|(a_{v})$, so $\mathrm{N}(\lambda)|\mathrm{N}((a_{v}))$ in $\mathbb{Z}$. By the assumption on $(a_{v})$,
$\mathrm{N}((a_{v}))=\mathrm{N}_{K/\mathbb{Q}}(a_{v})=\bigg{|}\prod_{\sigma:K\hookrightarrow\overline{\mathbb{Q}}}\sigma(a_{v})\bigg{|}=\prod_{\iota:K\rightarrow\mathbb{C}}|\iota(a_{v})|_{\mathbb{C}}\leq 2^{g}\mathrm{N}(v)^{g/2},$
where we write $[K:\mathbb{Q}]=g$. By the assumptions on $v$ and $p$, we have $\mathrm{N}(v)=p$ and $\mathrm{N}(\lambda)=p^{g}$. Hence, $p^{g}|\mathrm{N}((a_{v}))$ with $\mathrm{N}((a_{v}))\leq 2^{g}p^{g/2}$. As $p\geq 5$, this implies that $\mathrm{N}((a_{v}))=0$ so $a_{v}=0$. ∎
We continue with the proof of Theorem 1.7. Let $T_{1}$ be the set of rational primes $p$ such that
  1. (1)
    $p$ lies below a degree $1$ prime of $F$;
  2. (2)
    $p$ is inert in $K$;
  3. (3)
    $p\geq 5$.
  4. (4)
    $p$ does not lie below any prime in $S$.
By the assumption in Theorem 1.7, there exists a rational prime $p$ satisfying both (1) and (2), and by Proposition 2.7, there exists positive density rational primes having the same splitting type as $p$ in both $F$ and $K$. In particular, these primes all satisfy both (1) and (2), implying that $T_{1}$ has positive density. Now, consider the set of primes of $F$ that lie above the primes in $T_{1}$, i.e., consider
$S_{1}:=\{v\in\Sigma_{F}:\mathrm{N}(v)=p\text{ for some }p\in T_{1}\}.$
By Lemma 4.1, if $v\in S_{1}$, then $a_{v}=0$ or $v\in S^{\mathrm{ord}}$. Since $\{v\in\Sigma_{F}:v\notin S\text{ and }a_{v}=0\}$ is assumed to have density 00, $S_{1}$ thus has the same density as $S_{1}\cap\{v\in\Sigma_{F}:v\notin S\text{ and }a_{v}\neq 0\}\subseteq S^{\mathrm{ord}}$, so it suffices to prove that $S_{1}$ has positive density.
Note that each prime $p\in T_{1}$ gives at least one prime $v\in S_{1}$, so for $X>0$
#$\#\{v\in S_{1}:\mathrm{N}(v)\leq X\}\geq\#\{p\in T_{1}:p\leq X\}.$
On the other hand, by Landau’s prime ideal theorem (and the usual prime number theorem),
#$\#\{v\in\Sigma_{F}:\mathrm{N}(v)\leq X\}\sim\frac{X}{\log X}\sim\#\{p:p\leq X\}.$
Hence,
$\mathrm{Density}(S_{1})=\lim_{X\rightarrow\infty}\frac{\#\{v\in S_{1}:\mathrm{N}(v)\leq X\}}{\#\{v\in\Sigma_{F}:\mathrm{N}(v)\leq X\}}$ $\geq\lim_{X\rightarrow\infty}\frac{\#\{p\in T_{1}:p\leq X\}}{\#\{v\in\Sigma_{F}:\mathrm{N}(v)\leq X\}}$
$=\lim_{X\rightarrow\infty}\frac{\#\{p\in T_{1}:p\leq X\}}{\#\{p:p\leq X\}}$
$=\mathrm{Density}(T_{1})>0.$
Theorem 1.7 then follows.
5. Proof of Theorem 1.6
We will now give a proof of Theorem 1.6. We will adopt all the notations in Theorem 1.6 unless stated otherwise. Recall that in this case, we have a rational prime $p$ that lies below a degree $1$ prime of $F$ and splits into two distinct primes in $K$ of the same inertia degree.
5.1. A proof assuming Corollary 5.7
The proof of Theorem 1.6 follows a similar approach to that of Theorem 1.7. We first prove a lemma showing that under the assumption on the rational prime $p$, almost every prime $v$ of $F$ above $p$ is either an ordinary prime or belongs to a special set. In the case of Theorem 1.7, the special set is $\{v\in\Sigma_{F}:v\notin S\text{ and }a_{v}=0\}$, which has density 0 by assumptions. For Theorem 1.6, the special set is S(c) (see Equation 4 for its precise definition) or, more accurately, a finite union of such sets. It is much trickier to show that the density of $S(c)$ is 00. Therefore, we first proceed with the proof assuming that $S(c)$ has density 00, and then, in Section 5.2, we discuss its density properly.
Lemma 5.1.
Let $F$ and $K$ be two number fields and let $S$ be a finite set of primes of $F$. Let $(a_{v})_{v\in\Sigma_{F}\setminus S}$ be a fixed sequence in $\mathcal{O}_{K}$ satisfying that
$|\iota(a_{v})|_{\mathbb{C}}\leq 2\mathrm{N}(v)^{1/2}$
for all $v\in\Sigma_{F}$ with $v\notin S$ and all embeddings $\iota:K\rightarrow\mathbb{C}$. Let $v\in\Sigma_{F}$ with $v\notin S$ be a degree $1$ prime of $F$ and let $p$ be the rational prime lying below $v$. Suppose that $p$ splits into two distinct primes in $K$ of the same inertia degree. Then, $v\in S^{\mathrm{ord}}$ or $\mathrm{N}_{K/\mathbb{Q}}(a_{v})^{2}=c\mathrm{N}(v)^{g}$ for some $c\in\mathbb{Z}$ with $0\leq c\leq 2^{2g}$.
Proof.
Suppose that $v\notin S^{\mathrm{ord}}$. Then, there exists a prime $\lambda$ of $K$ with $\lambda|\mathrm{N}(v)$ such that $\lambda|(a_{v})$, so $\mathrm{N}(\lambda)|\mathrm{N}((a_{v}))$ in $\mathbb{Z}$. By the assumption on $a_{v}$,
$\mathrm{N}((a_{v}))=\bigg{|}\prod_{\sigma:K\hookrightarrow\overline{\mathbb{Q}}}\sigma(a_{v})\bigg{|}=\prod_{\iota:K\rightarrow\mathbb{C}}|\iota(a_{v})|_{\mathbb{C}}\leq 2^{g}\mathrm{N}(v)^{g/2},$
where we write $[K:\mathbb{Q}]=g$. By the assumptions on $v$ and $p$, we have $\mathrm{N}(v)=p$ and $\mathrm{N}(\lambda)=p^{g/2}$ ($g$ is necessarily even in this case). Hence, $p^{g/2}|\mathrm{N}((a_{v}))$ with $\mathrm{N}((a_{v}))\leq 2^{g}p^{g/2}$, implying that $\mathrm{N}_{K/\mathbb{Q}}(a_{v})^{2}=c\mathrm{N}(v)^{g}$ for some $c\in\mathbb{Z}$ with $0\leq c\leq 2^{2g}$. ∎
Now, we continue with the proof. For $c\in\mathbb{Z}$, let
$S(c):=\{v\in\Sigma_{F}:v\notin S\text{ and }\mathrm{N}_{K/\mathbb{Q}}(a_{v})^{2}=c\mathrm{N}(v)^{g}\}.$ (4)
Suppose for now that $S(c)$ has density 00, which will be proved in Corollary 5.7. We will show under this assumption that Theorem 1.6 holds. Let $T_{2}$ be the set of rational primes $p$ such that
  1. (1)
    $p$ lies below a degree $1$ prime of $F$;
  2. (2)
    $p$ splits into two primes in $K$ of the same inertia degree.
  3. (3)
    $p$ does not lie below any prime in $S$.
As before, $T_{2}$ has positive density by Proposition 2.7 and the assumption on the existence of a rational prime satisfying both (1) and (2). Let
$S_{2}:=\{v\in\Sigma_{F}:\mathrm{N}(v)=p\text{ for some }p\in T_{2}\}.$
By Lemma 5.1, if $v\in S_{2}$, then $v\in S^{\mathrm{ord}}$ or $v\in S(c)$ for some $c\in\mathbb{Z}$ with $0\leq c\leq 2^{2g}$. As $S(c)$ and hence $\bigcup_{0\leq c\leq 2^{2g}}S(c)$ have density 00, it follows that $S_{2}$ has the same density as $S_{2}\cap S^{\mathrm{ord}}\subseteq S^{\mathrm{ord}}$, so it suffices to prove that $S_{2}$ has positive density. Now, the same argument as before gives that
$\mathrm{Density}(S_{2})=\lim_{X\rightarrow\infty}\frac{\#\{v\in S_{2}:\mathrm{N}(v)\leq X\}}{\#\{v\in\Sigma_{F}:\mathrm{N}(v)\leq X\}}\geq\lim_{X\rightarrow\infty}\frac{\#\{p\in T_{2}:p\leq X\}}{\#\{p:p\leq X\}}=\mathrm{Density}(T_{2})>0.$
Thus, Theorem 1.6 follows from the argument above.
5.2. Density of $S(c)$
The main goal of this subsection is to show that $S(c)$ has density 00 for each $c\in\mathbb{Z}$ with $0\leq c\leq 2^{2g}$, so that Theorem 1.6 holds by Section 5.1. To prove this, we need to use the surjective property of the family of Galois representations $(\rho_{\ell})_{\ell\in T}$.
Fix a rational prime $\ell$ that splits completely in $K$ and fix an embedding $\iota_{\ell}:\overline{\mathbb{Q}}\hookrightarrow\overline{\mathbb{Q}_{\ell}}$. Write $[K^{\prime}:\mathbb{Q}]=g^{\prime}$ and let $\sigma^{\prime}_{1},\ldots,\sigma^{\prime}_{g^{\prime}}:K^{\prime}\hookrightarrow\overline{\mathbb{Q}}$ be all the embeddings of $K^{\prime}$ into $\overline{\mathbb{Q}}$. As $\ell$ splits completely in $K$ and hence in $K^{\prime}$, by Lemma 2.2, $\{\lambda^{\prime}_{i}:=\lambda^{\prime}_{\sigma^{\prime}_{i}}\mid 1\leq i\leq g^{\prime}\}$ is an enumeration of all the places $\lambda^{\prime}|\ell$ of $K^{\prime}$. In particular, we may simply make the identification $K^{\prime}_{\lambda^{\prime}_{i}}=\widehat{\iota_{\ell}\sigma^{\prime}_{i}(K^{\prime})}=\mathbb{Q}_{\ell}\subseteq\overline{\mathbb{Q}_{\ell}}$ so that the natural map $K^{\prime}\rightarrow K^{\prime}_{\lambda^{\prime}_{i}}$ is given by $\alpha\mapsto\iota_{\ell}\sigma^{\prime}_{i}(\alpha)$. In this way,
$\mathcal{O}_{K^{\prime}}\otimes\mathbb{Z}_{\ell}$ $\xrightarrow{\;\sim\;}\prod_{i=1}^{g}\mathcal{O}_{K^{\prime}_{\lambda^{\prime}_{i}}}$
$\alpha\otimes\beta$ $\mapsto(\iota_{\ell}\sigma^{\prime}_{i}(\alpha)\cdot\beta)_{i=1}^{g^{\prime}}.$
Write $[K:K^{\prime}]=d$. Since $\ell$ splits completely in $K$, each $\lambda^{\prime}_{i}$ also splits completely in $K$. In particular, for every prime $\lambda|\lambda^{\prime}_{i}$ of $K$, we have $\mathcal{O}_{K^{\prime}_{\lambda^{\prime}_{i}}}=\mathcal{O}_{K_{\lambda}}$. Let us now enumerate the primes $\lambda|\ell$ of $K$ by $(\lambda_{ij})_{i=1,j=1}^{g^{\prime},\;d}$ such that $\prod_{j=1}^{d}\lambda_{ij}=\lambda^{\prime}_{i}$ and let $\sigma_{ij}:K\hookrightarrow\overline{\mathbb{Q}}$ be the embedding corresponding to the prime $\lambda_{ij}$ by Lemma 2.2. We also make the identification $K_{\lambda_{ij}}=\widehat{\iota_{\ell}\sigma_{ij}(K)}=\mathbb{Q}_{\ell}\subseteq\overline{\mathbb{Q}_{\ell}}$ so that similarly
$\mathcal{O}_{K}\otimes\mathbb{Z}_{\ell}$ $\xrightarrow{\;\sim\;}\prod_{i=1}^{g^{\prime}}\prod_{j=1}^{d}\mathcal{O}_{K_{\lambda_{ij}}}$
$\alpha\otimes\beta$ $\mapsto(\iota_{\ell}\sigma_{ij}(\alpha)\cdot\beta)_{i=1,j=1}^{g^{\prime},\;d}.$
Thus, the composition map
$\prod_{i=1}^{g^{\prime}}\mathcal{O}_{K^{\prime}_{\lambda^{\prime}_{i}}}\xrightarrow{\;\sim\;}\mathcal{O}_{K^{\prime}}\otimes\mathbb{Z}_{\ell}\hookrightarrow\mathcal{O}_{K}\otimes\mathbb{Z}_{\ell}\xrightarrow{\;\sim\;}\prod_{i=1}^{g^{\prime}}\prod_{j=1}^{d}\mathcal{O}_{K_{\lambda_{ij}}}$
is given by
$\prod_{i=1}^{g^{\prime}}\mathcal{O}_{K^{\prime}_{\lambda^{\prime}_{i}}}$ $\hookrightarrow\prod_{i=1}^{g^{\prime}}\prod_{j=1}^{d}\mathcal{O}_{K_{\lambda_{ij}}}$
$(a_{i})_{1\leq i\leq g^{\prime}}$ $\mapsto((a_{i})_{1\leq j\leq d})_{1\leq i\leq g^{\prime}}.$
In this way, we obtain an isomorphism
$\operatorname{GL}_{2}(\mathcal{O}_{K}\otimes\mathbb{Z}_{\ell})\xrightarrow{\;\sim\;}\prod_{i=1}^{g^{\prime}}\prod_{j=1}^{d}\operatorname{GL}_{2}(\mathcal{O}_{K_{\lambda_{ij}}})\simeq\prod_{i=1}^{g^{\prime}}\prod_{j=1}^{d}\operatorname{GL}_{2}(\mathbb{Z}_{\ell}),$ (5)
and $\operatorname{GL}_{2}(\mathcal{O}_{K^{\prime}}\otimes\mathbb{Z}_{\ell})$ is identified with $\prod_{i=1}^{g^{\prime}}\Delta_{d}(\operatorname{GL}_{2}(\mathbb{Z}_{\ell}))$ under this isomorphism, where $\Delta_{d}:\operatorname{GL}_{2}(\mathbb{Z}_{\ell})\rightarrow\prod_{j=1}^{d}\operatorname{GL}_{2}(\mathbb{Z}_{\ell})$ is the diagonal embedding.
Now, let $\ell\in T$ (which is assumed to split completely in $K$). Let $\rho^{\prime}_{\ell}:\operatorname{Gal}(\overline{F}/F)\rightarrow\prod_{i=1}^{g^{\prime}}\prod_{j=1}^{d}\operatorname{GL}_{2}(\mathbb{Z}_{\ell})$ be the composition of $\rho_{\ell}$ and the isomorphism (5) and let $\overline{\rho^{\prime}_{\ell}}:\operatorname{Gal}(\overline{F}/F)\rightarrow\prod_{i=1}^{g^{\prime}}\prod_{j=1}^{d}\operatorname{GL}_{2}(\mathbb{F}_{\ell})$ be the composition of $\rho^{\prime}_{\ell}$ and the natural reduction map $\mathbb{Z}_{\ell}\rightarrow\mathbb{F}_{\ell}$. Recall that for all primes $\ell\in T$,
$\rho_{\ell}(\operatorname{Gal}(\overline{F}/F^{\prime}))=\{M\in\operatorname{GL}_{2}(\mathcal{O}_{K^{\prime}}\otimes\mathbb{Z}_{\ell}):\operatorname{det}M\in(\mathbb{Z}\otimes\mathbb{Z}_{\ell})^{\times}(\simeq\mathbb{Z}_{\ell}^{\times})\}.$
From this, we obtain that
[…]
Now, since $[F^{\prime}:F]$ is finite, there exists $n\in\mathbb{Z}^{+}$ such that $\tau^{n}\in\operatorname{Gal}(\overline{F}/F^{\prime})$ for all $\tau\in\operatorname{Gal}(\overline{F}/F)$. [^8] By considering all the cosets $\{\tau^{i}\operatorname{Gal}(\overline{F}/F^{\prime})\}_{i\in\mathbb{Z}}$, one can deduce that $\tau^{i}\in\operatorname{Gal}(\overline{F}/F^{\prime})$ for some $1\leq i\leq[F^{\prime}:F]$, so $n$ can be chosen as $[F^{\prime}:F]!$. In particular, this means that
$\operatorname{Im}\overline{\rho^{\prime}_{\ell}}\subseteq\widetilde{B}_{\ell}(g^{\prime},d,n):=$$\big{\{}M\in\prod_{i=1}^{g^{\prime}}\prod_{j=1}^{d}\operatorname{GL}_{2}(\mathbb{F}_{\ell}):M^{n}\in A_{\ell}(g^{\prime},d)\big{\}}$
$=$ $\big{\{}(M_{ij})_{i=1,j=1}^{g^{\prime},\;d}\in\prod_{i=1}^{g^{\prime}}\prod_{j=1}^{d}\operatorname{GL}_{2}(\mathbb{F}_{\ell}):M_{ij}^{n}=M_{ij^{\prime}}^{n}\text{ for all }1\leq j,j^{\prime}\leq d,$
[…]
In fact, $\operatorname{Im}\overline{\rho^{\prime}_{\ell}}$ lies in a smaller set when $\ell>n+1$. To prove this, we first need a lemma on the matrix group $\operatorname{GL}_{2}(\mathbb{F}_{\ell})$.
Lemma 5.2.
Let $n\in\mathbb{Z}^{+}$ and let $\ell>n+1$ be a prime. Let $t\in\mathbb{F}_{\ell}^{\times}$ such that $t^{n}\neq 1$. Then, $M\in\operatorname{GL}_{2}(\mathbb{F}_{\ell})$ commutes with both $\begin{pmatrix}1&0\\ 0&t^{n}\end{pmatrix}$ and $\begin{pmatrix}1&\overline{n}\\ 0&1\end{pmatrix}$ if and only if $M=\begin{pmatrix}\lambda&0\\ 0&\lambda\end{pmatrix}$ for some $\lambda\in\mathbb{F}_{\ell}^{\times}$, where $\overline{n}$ denotes $n\pmod{\ell}$.
Proof.
Let $M=\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\operatorname{GL}_{2}(\mathbb{F}_{\ell})$. Then, $M\begin{pmatrix}1&0\\ 0&t^{n}\end{pmatrix}=\begin{pmatrix}1&0\\ 0&t^{n}\end{pmatrix}M$ implies that
$\begin{pmatrix}a&t^{n}b\\ c&t^{n}d\end{pmatrix}=\begin{pmatrix}a&b\\ t^{n}c&t^{n}d\end{pmatrix}.$
Hence, $b(t^{n}-1)=c(t^{n}-1)=0$, implying that $b=c=0$ as $t^{n}\neq 1$. Now, $M\begin{pmatrix}1&\overline{n}\\ 0&1\end{pmatrix}=\begin{pmatrix}1&\overline{n}\\ 0&1\end{pmatrix}M$ implies that
$\begin{pmatrix}a&a\overline{n}\\ 0&d\end{pmatrix}=\begin{pmatrix}a&d\overline{n}\\ 0&d\end{pmatrix}.$
Hence, $a\overline{n}=d\overline{n}$, implying that $a=d$ as $\overline{n}\in\mathbb{F}_{\ell}^{\times}$. ∎
Proposition 5.3.
Let $n\in\mathbb{Z}^{+}$ be such that $\tau^{n}\in\operatorname{Gal}(\overline{F}/F^{\prime})$ for all $\tau\in\operatorname{Gal}(\overline{F}/F)$ and let $\ell\in T$ be a prime with $\ell>n+1$. Then,
$\operatorname{Im}\overline{\rho^{\prime}_{\ell}}\subseteq B_{\ell}(g^{\prime},d,n):=$$\big{\{}(M_{ij})_{i=1,j=1}^{g^{\prime},\;d}\in\prod_{i=1}^{g^{\prime}}\prod_{j=1}^{d}\operatorname{GL}_{2}(\mathbb{F}_{\ell}):M_{i1}^{-1}M_{ij}\in\mu_{n}(\mathbb{F}_{\ell})I_{2},$
[…]
where $\mu_{n}(\mathbb{F}_{\ell}):=\{\zeta\in\mathbb{F}_{\ell}:\zeta^{n}=1\}$.
Remark 5.4.
Under the notation in the proposition, there is a natural identification
$B_{\ell}(g^{\prime},1,n)\times\mu_{n}(\mathbb{F}_{\ell})^{g^{\prime}(d-1)}=B_{\ell}(g^{\prime},1,n)\times\prod_{i=1}^{g^{\prime}}\prod_{j=2}^{d}\mu_{n}(\mathbb{F}_{\ell})$ $\xrightarrow{\;\sim\;}B_{\ell}(g^{\prime},d,n)$
$\big{(}(M_{i})_{i=1}^{g^{\prime}},(\zeta_{ij})_{i=1,j=2}^{g^{\prime},\;d}\big{)}$ $\longmapsto(\zeta_{ij}M_{i})_{i=1,j=1}^{g^{\prime},\;d},$
where $\zeta_{i1}:=1$ for all $1\leq i\leq g^{\prime}$.
Proof.
Fix $\tau\in\operatorname{Gal}(\overline{F}/F)$ and write $\overline{\rho^{\prime}_{\ell}}(\tau)=M=(M_{ij})_{i=1,j=1}^{g^{\prime},\;d}$. Let $A=(A_{i})_{i=1}^{g^{\prime}}\in A_{\ell}(g^{\prime},d)=\overline{\rho^{\prime}_{\ell}}(\operatorname{Gal}(\overline{F}/F^{\prime}))$ and let $\alpha\in\operatorname{Gal}(\overline{F}/F^{\prime})$ be such that $\overline{\rho^{\prime}_{\ell}}(\alpha)=A$. Then,
$MAM^{-1}=\overline{\rho^{\prime}_{\ell}}(\tau\alpha\tau^{-1})\in\widetilde{B}_{\ell}(g^{\prime},d,n).$
In particular,
$M_{ij}A_{i}^{n}M_{ij}^{-1}=(M_{ij}A_{i}M_{ij}^{-1})^{n}=(M_{i1}A_{i}M_{i1}^{-1})^{n}=M_{i1}A_{i}^{n}M_{i1}^{-1}$
or equivalently
$(M_{i1}^{-1}M_{ij})A_{i}^{n}=A_{i}^{n}(M_{i1}^{-1}M_{ij})$
for all $1\leq i\leq g^{\prime}$ and $1\leq j\leq d$.
Since $A\in A_{\ell}(g^{\prime},d)$ is arbitrary and $\ell>n+1$, we may pick $A$ with $A_{i}$ being $\begin{pmatrix}1&0\\ 0&t\end{pmatrix}$ or $\begin{pmatrix}1&1\\ 0&1\end{pmatrix}$, where $t\in\mathbb{F}_{\ell}^{\times}$ with $t^{n}\neq 1$, so that $A_{i}^{n}$ is $\begin{pmatrix}1&0\\ 0&t^{n}\end{pmatrix}$ or $\begin{pmatrix}1&n\\ 0&1\end{pmatrix}$. In particular, it follows from Lemma 5.2 that
$M_{i1}^{-1}M_{ij}=\begin{pmatrix}\lambda&0\\ 0&\lambda\end{pmatrix}$
for some $\lambda\in\mathbb{F}_{\ell}^{\times}$. To see that $\lambda\in\mu_{n}(\mathbb{F}_{\ell})$, note that $(M_{ij})_{i=1,j=1}^{g^{\prime},\;d}=\overline{\rho^{\prime}_{\ell}}(\tau)\in\widetilde{B}_{\ell}(g^{\prime},d,n)$, so
$M_{ij}^{n}=M_{i^{\prime}j^{\prime}}^{n}\text{ for all }1\leq i,i^{\prime}\leq g^{\prime}\text{ and }1\leq j,j^{\prime}\leq d.$
In particular,
$\begin{pmatrix}\lambda^{n}&0\\ 0&\lambda^{n}\end{pmatrix}=(M_{i1}^{-1}M_{ij})^{n}=M_{i1}^{-n}M_{ij}^{n}=I_{2}$
(note that $M_{ij}$ and $M_{i1}$ commute). Hence, $\lambda\in\mu_{n}(\mathbb{F}_{\ell})$. ∎
Let $v\notin S$ be a prime of $F$ and fix a Frobenius lift $\operatorname{Frob}_{v}\in\operatorname{Gal}(\overline{F}/F)$ at $v$. Let $\rho^{\prime}_{\ell}(\operatorname{Frob}_{v})=(A_{ij})_{i=1,j=1}^{g^{\prime},\;d}$. Then, $\operatorname{tr}A_{ij}=\iota_{\ell}\sigma_{ij}(a_{v})$ and $\operatorname{det}A_{ij}=\iota_{\ell}\sigma_{ij}\big{(}\epsilon(v)\mathrm{N}(v)\big{)}$ for all $1\leq i\leq g^{\prime}$ and $1\leq j\leq d$. Note that $\epsilon$ is of finite order, so $\mathrm{N}_{K/\mathbb{Q}}(\epsilon(v))=\pm 1$, and hence $\mathrm{N}_{K/\mathbb{Q}}(\epsilon(v))^{g}=1$ (as $g$ is necessarily even). In this way, the condition that $\mathrm{N}_{K/\mathbb{Q}}(a_{v})^{2}=c\mathrm{N}(v)^{g}$ is equivalent to that
$\bigg{(}\prod_{i=1}^{g^{\prime}}\prod_{j=1}^{d}\operatorname{tr}A_{ij}\bigg{)}^{2}=c\cdot\prod_{i=1}^{g^{\prime}}\prod_{j=1}^{d}\operatorname{det}A_{ij}.$ (6)
As this condition stays the same under the reduction map $\mathbb{Z}_{\ell}\rightarrow\mathbb{F}_{\ell}$, we obtain that for $v\in S(c)$,
$\overline{\rho^{\prime}_{\ell}}(\operatorname{Frob}_{v})\in C_{\ell}(g^{\prime},d,n,c)\cap\operatorname{Im}\overline{\rho^{\prime}_{\ell}}=\bigg{\{}(M_{ij})_{i=1,j=1}^{g^{\prime},\;d}\in\operatorname{Im}\overline{\rho^{\prime}_{\ell}}:\bigg{(}\prod_{i=1}^{g^{\prime}}\prod_{j=1}^{d}\operatorname{tr}M_{ij}\bigg{)}^{2}=\bar{c}\cdot\prod_{i=1}^{g^{\prime}}\prod_{j=1}^{d}\operatorname{det}M_{ij}\bigg{\}},$
where $\bar{c}$ denotes $c\pmod{\ell}$ and
[…]
Note that $C_{\ell}(g^{\prime},d,n,c)\cap\operatorname{Im}\overline{\rho^{\prime}_{\ell}}$ is invariant under conjugation by $\operatorname{Im}\overline{\rho^{\prime}_{\ell}}$. By the Chebotarev density theorem,
$\mathrm{Density}(S(c))\leq\frac{|C_{\ell}(g^{\prime},d,n,c)\cap\operatorname{Im}\overline{\rho^{\prime}_{\ell}}|}{|\operatorname{Im}\overline{\rho^{\prime}_{\ell}}|}\leq\frac{|C_{\ell}(g^{\prime},d,n,c)|}{|A_{\ell}(g^{\prime},d)|}.$ (7)
We will now prove that the latter expression goes to 00 as $\ell\to\infty$, from which it follows that $S(c)$ has density 00.
Lemma 5.5.
Let $\ell$ be an odd prime, $t\in\mathbb{F}_{\ell}$, and $d\in\mathbb{F}_{\ell}^{\times}$. Then we have
#$\#\left\{N\in\operatorname{GL}_{2}(\mathbb{Z}/\ell\mathbb{Z}):\operatorname{det}N=d,\operatorname{tr}N=t\right\}=\ell^{2}+\ell\cdot\bigg{(}\frac{t^{2}-4d}{\ell}\bigg{)},$
where $\left(\frac{\cdot}{\ell}\right)$ is the Legendre symbol.
Proof.
See [CFM05, Lemma 2.7]. ∎
Lemma 5.6.
Let $c_{0}\in\mathbb{Z}$. Then, for all sufficiently large prime $\ell$, we have
$\frac{|C_{\ell}(g^{\prime},d,n,c_{0})|}{|A_{\ell}(g^{\prime},d)|}\ll_{g^{\prime},d,n}\frac{1}{\ell}.$
Proof.
First, we give an asymptotic of $|A_{\ell}(g^{\prime},d)|$. Observe that
$|A_{\ell}(g^{\prime},d)|=|A_{\ell}(g^{\prime})|,$
where for any $g\geq 1$,
$A_{\ell}(g):=\bigg{\{}(N_{i})_{1\leq i\leq g}\in\prod_{1\leq i\leq g}\operatorname{GL}_{2}(\mathbb{F}_{\ell}):\operatorname{det}N_{i}=\operatorname{det}N_{i^{\prime}}\text{ for all $1\leq i,i^{\prime}\leq g$}\bigg{\}}.$
It has been computed in [CW23, Lemma 12] that $|A_{\ell}(g^{\prime})|\asymp\ell^{3g^{\prime}+1}.$ Hence, $|A_{\ell}(g^{\prime},d)|\asymp\ell^{3g^{\prime}+1}.$
Next, we give an upper bound of $|C_{\ell}(g^{\prime},d,n,c_{0})|$. By Proposition 5.3, we observe that
$|C_{\ell}(g^{\prime},d,n,c_{0})|\leq n^{g^{\prime}d}\sum_{c\in\mu_{n}(\mathbb{F}_{\ell})}|D_{\ell}(g^{\prime},d,n,c{\hskip 1.0pt}\overline{c_{0}})|,$ (8)
where for any $c\in\mathbb{F}_{\ell}$, we define
$D_{\ell}(g^{\prime},d,n,c)$ :=$:=\bigg{\{}(N_{i})_{1\leq i\leq g^{\prime}}\in\prod_{1\leq i\leq g^{\prime}}\operatorname{GL}_{2}(\mathbb{F}_{\ell}):(\operatorname{det}N_{i})^{n}=(\operatorname{det}N_{i^{\prime}})^{n}\text{ for all $1\leq i,i^{\prime}\leq g^{\prime}$}$
$\hskip 85.35826pt\text{and }\bigg{(}\prod_{1\leq i\leq g^{\prime}}\operatorname{tr}N_{i}\bigg{)}^{2d}=c\cdot\prod_{1\leq i\leq g^{\prime}}\left(\operatorname{det}N_{i}\right)^{d}\bigg{\}}.$
To see this, for any $(N_{i,j})_{i=1,j=1}^{g^{\prime},\;d}\in C_{\ell}(g^{\prime},d,n,c_{0})$, we have $N_{i,j}=\zeta_{i,j}N_{i,1}$ with $\zeta_{i,j}\in\mu_{n}(\mathbb{F}_{\ell})$ for all $1\leq i\leq g^{\prime}$ and $1\leq j^{\prime}\leq d$. By denoting $N_{i}:=N_{i,1}$, we see that there exist $c_{1},c_{2}\in\mu_{n}(\mathbb{F}_{\ell})$ such that
$\bigg{(}\prod_{1\leq i\leq g^{\prime}}\prod_{1\leq j\leq d}\operatorname{tr}N_{i,j}\bigg{)}^{2}=c_{1}\bigg{(}\prod_{1\leq i\leq g^{\prime}}\operatorname{tr}N_{i}\bigg{)}^{2d}\;\text{ and }\;\prod_{1\leq i\leq g^{\prime}}\prod_{1\leq j\leq d}\operatorname{det}N_{i,j}=c_{2}\bigg{(}\prod_{1\leq i\leq g^{\prime}}\operatorname{det}N_{i}\bigg{)}^{d}.$
Finally, the factor $n^{g^{\prime}d}$ in Equation 8 arises from the possible choices of $\zeta_{i,j}$. Hence, it suffices to estimate $|D_{\ell}(g^{\prime},d,n,c{\hskip 1.0pt}\overline{c_{0}})|$ for $c\in\mu_{n}(\mathbb{F}_{\ell})$. Observe that if $g^{\prime}=1$, then by Lemma 5.5,
$|D_{\ell}(1,d,n,c{\hskip 1.0pt}\overline{c_{0}})|$ =#$=\#\big{\{}N\in\operatorname{GL}_{2}(\mathbb{F}_{\ell}):(\operatorname{tr}N)^{2d}=c{\hskip 1.0pt}\overline{c_{0}}\cdot(\operatorname{det}N)^{d}\big{\}}$
$\leq\sum_{D\in\mathbb{F}_{\ell}^{\times}}\#\{N\in\operatorname{GL}_{2}(\mathbb{F}_{\ell}):(\operatorname{tr}N)^{2d}=c{\hskip 1.0pt}\overline{c_{0}}D^{d},\operatorname{det}N=D\}$
$\ll(\ell-1)\cdot(2d)\ell^{2}\ll_{d}\ell^{3}(=\ell^{3g^{\prime}}).$
For $g^{\prime}\geq 2$, we will compare the size of $D_{\ell}(g^{\prime},d,n,c{\hskip 1.0pt}\overline{c_{0}})$ with the following matrix group
$A_{\ell,n}(g^{\prime}):=\bigg{\{}(N_{i})_{1\leq i\leq g^{\prime}}\in\prod_{1\leq i\leq g^{\prime}}\operatorname{GL}_{2}(\mathbb{F}_{\ell}):(\operatorname{det}N_{i})^{n}=(\operatorname{det}N_{i^{\prime}})^{n}\text{ for all $1\leq i,i^{\prime}\leq g^{\prime}$}\bigg{\}}.$
Since for each $1\leq i,i^{\prime}\leq g^{\prime}$, $(\operatorname{det}N_{i})^{n}=(\operatorname{det}N_{i^{\prime}})^{n}$ implies $\operatorname{det}N_{i}\in\mu_{n}(\mathbb{F}_{\ell})\operatorname{det}N_{i}^{\prime}$, we obtain
$|A_{\ell,n}(g^{\prime})|\leq n^{g^{\prime}}|A_{\ell}(g^{\prime})|\asymp_{n,g^{\prime}}\ell^{3g^{\prime}+1}.$
Therefore, if $\overline{c_{0}}\neq 0$, then
$|D_{\ell}(g^{\prime},d,n,c{\hskip 1.0pt}\overline{c_{0}})|$ $\leq\sum_{(N_{i})_{1\leq i\leq g^{\prime}-1}\in A_{\ell,n}(g^{\prime}-1)}\#\bigg{\{}N_{g^{\prime}}\in\operatorname{GL}_{2}(\mathbb{F}_{\ell}):\operatorname{det}N_{1}^{n}=\dots=\operatorname{det}N_{g^{\prime}}^{n},$
$\hskip 142.26378pt\prod_{1\leq i\leq g^{\prime}-1}(\operatorname{tr}N_{i})^{2d}\cdot(\operatorname{tr}N_{g^{\prime}})^{2d}=c{\hskip 1.0pt}\overline{c_{0}}\cdot(\operatorname{det}N_{g^{\prime}})^{d}\bigg{\}}$
$=\sum_{(N_{i})_{1\leq i\leq g^{\prime}-1}\in A_{\ell,n}(g^{\prime}-1)}\#\bigg{\{}N_{g^{\prime}}\in\operatorname{GL}_{2}(\mathbb{F}_{\ell}):\operatorname{det}N_{g^{\prime}}^{n}=D,\;T^{n}\cdot(\operatorname{tr}N_{g^{\prime}})^{2dn}=c{\hskip 1.0pt}\overline{c_{0}}D^{d},$
$\hskip 156.49014pt\text{where $D:=\operatorname{det}N_{1}^{n}$, $T:=\prod_{1\leq i\leq g^{\prime}-1}(\operatorname{tr}N_{i})^{2d}(\neq 0)$}\bigg{\}}$
$\ll|A_{\ell,n}(g^{\prime}-1)|\cdot(2dn^{2})\ell^{2}$
$\ll_{d,n}\ell^{3g^{\prime}}.$
If $\overline{c_{0}}=0$, then we get
$|D_{\ell}(g^{\prime},d,n,0)|$ $\leq\sum_{(N_{i})_{1\leq i\leq g^{\prime}-1}\in D_{\ell}(g^{\prime}-1,d,n,0)}\#\bigg{\{}N_{g^{\prime}}\in\operatorname{GL}_{2}(\mathbb{F}_{\ell}):\operatorname{det}N_{1}^{n}=\dots=\operatorname{det}N_{g^{\prime}}^{n}\bigg{\}}$
$\hskip 28.45274pt+\sum_{(N_{i})_{1\leq i\leq g^{\prime}-1}\in A_{\ell,n}(g^{\prime}-1)}\#\bigg{\{}N_{g^{\prime}}\in\operatorname{GL}_{2}(\mathbb{F}_{\ell}):\operatorname{det}N_{g^{\prime}}^{n}=D,\;T\cdot(\operatorname{tr}N_{g^{\prime}})^{2d}=0,$
$\hskip 159.3356pt\text{where $D:=\operatorname{det}N_{1}^{n},\;T:=\prod_{1\leq i\leq g^{\prime}-1}(\operatorname{tr}N_{i})^{2d}(\neq 0)$}\bigg{\}}$
$\ll|D_{\ell}(g^{\prime}-1,d,n,0)|\cdot n\ell^{3}+|A_{\ell,n}(g^{\prime}-1)|\cdot n\ell^{2}$
$\ll_{d,n}\ell^{3g^{\prime}},$
where in the last line we use the induction hypothesis $|D_{\ell}(g^{\prime}-1,d,n,0)|\ll\ell^{3(g^{\prime}-1)}$ on $g^{\prime}-1$. In all, for all sufficiently large $\ell$, we have
$\frac{|C_{\ell}(g^{\prime},d,n,cc_{0})|}{|A_{\ell}(g^{\prime},d)|}\ll_{g^{\prime},d,n}\frac{\ell^{3g^{\prime}}}{\ell^{3g^{\prime}+1}}=\frac{1}{\ell}.$
Corollary 5.7.
Under the assumptions on the family of Galois representations in Theorem 1.6, the set $S(c)$ has density 00.
Proof.
The proof follows easily from Equation 7 and Lemma 5.6. ∎
6. Applications of Theorem 1.6 and Theorem 1.7
6.1. Applications to abelian varieties
We now provide the proof of Theorem 1.1, assuming Theorem 1.5 (whose proof will be given in Section 6.2), the proof of Theorem 1.3 using Theorem 1.6 and Theorem 1.7, and the proof of Theorem 1.2 using Theorem 1.3. We adopt all the notations from Theorem 1.1, Theorem 1.2, and Theorem 1.3, respectively.
Proof of Theorem 1.1.
By the assumptions, $A$ is a $\operatorname{GL}_{2}$-type abelian variety. By the proof of [Rib04, Theorem 2.1] (see also [Wu20, Proposition 1.5]), $A$ is isogenous over $\mathbb{Q}$ to $B^{r}$ for some simple $\operatorname{GL}_{2}$-type abelian variety $B$ over $\mathbb{Q}$. Moreover, by [Rib04, Theorem 2.1] and its proof, $K^{\prime}:=\operatorname{End}_{\mathbb{Q}}(B)\otimes\mathbb{Q}$ is a number field of degree equal to $\dim B$ with $K^{\prime}\hookrightarrow K\hookrightarrow\operatorname{End}_{\mathbb{Q}}(A)\otimes\mathbb{Q}$. Since $[K:\mathbb{Q}]=q$ (resp. $2q$), $[K^{\prime}:\mathbb{Q}]=1$ or $q$ (resp. $1$, $2$, $q$, or $2q$). If $[K^{\prime}:\mathbb{Q}]=1$, then $B/\mathbb{Q}$ is an elliptic curve. In particular, the set of ordinary primes of $B$ has positive density, and so does the set of ordinary primes of $A$.
Suppose that $[K^{\prime}:\mathbb{Q}]\neq 1$. By [Rib04, Theorem 4.4], under Serre’s modularity conjecture, $B$ is isogenous over $\mathbb{Q}$ to abelian varieties of the form $A_{f}$ associated to some newform $f$ of weight $2$ under the Eichler–Shimura theory. As Serre’s modularity conjecture is proved in [KW09a, KW09b], we can thus complete the proof by invoking Theorem 1.5, Proposition 3.12, and the isomorphism $K^{\prime}=\operatorname{End}_{\mathbb{Q}}(A_{f})\otimes\mathbb{Q}\simeq K_{f}$. ∎
Proof of Theorem 1.3.
Recall that $A$ is absolutely simple, $\operatorname{End}_{F}(A)\otimes\mathbb{Q}$ contains a number field $K$ of degree $g$, and $A/\overline{F}$ is of the first kind. From these, it is easy to check that $A$ is of $\operatorname{GL}_{2}(K)$-type and satisfies the assumptions in Lemma 3.5 and Theorem 3.6. In particular, by the discussions in Section 3.1, the family of $\ell$-adic Galois representations $\{\rho_{A,\ell}:\operatorname{Gal}(\overline{F}/F)\rightarrow\operatorname{GL}_{2}(K\otimes\mathbb{Q}_{\ell})\}_{\ell\in\Sigma_{\mathbb{Q}}}$ satisfies that
  1. (1)
    each $\rho_{A,\ell}$ is unramified outside $\{v\in\Sigma_{F}:v|N_{A}\ell\}$;
  2. (2)
    for every primes $v\nmid N_{A}\ell$ of $F$, $\operatorname{det}(XI-\rho_{A,\ell}(\operatorname{Frob}_{v}))=X^{2}-(a_{v}(A)\otimes 1)X+(\epsilon(v)\mathrm{N}(v)\otimes 1)$, which is independent of $\ell$ (where $\epsilon:\operatorname{Gal}(\overline{F}/F)\rightarrow K^{\times}$ is some finite order character independent of $\ell$);
  3. (3)
    $|\iota(a_{v}(A))|_{\mathbb{C}}\leq 2\sqrt{\mathrm{N}(v)}$ for all primes $v\nmid N_{A}$ of $F$ and all embeddings $\iota:K\hookrightarrow\mathbb{C}$;
  4. (4)
    there exist two finite extensions of fields $F^{\prime}/F$ and $K/K^{\prime}$ such that for $\ell$ sufficiently large, we have, up to conjugation, that
    $\rho_{A,\ell}(\operatorname{Gal}(\overline{F^{\prime}}/F^{\prime}))=\{M\in\operatorname{GL}_{2}(\mathcal{O}_{K^{\prime}}\otimes\mathbb{Z}_{\ell}):\operatorname{det}M\in\mathbb{Z}_{\ell}^{\times}\}.$
Therefore, if we are in the case of (P2), then the result readily follows from Theorem 1.6 (with $T$ being the set of rational primes that split completely in $K$) and Proposition 3.8. In the case of (P1), we need to show that the set $\{v\in\Sigma_{F}:v\nmid N_{A}\text{ and }a_{v}(A)=0\}$ has density 00, which simply follows from Corollary 5.7 when $c=0$. The case of (P1) then follows from Theorem 1.7 and Proposition 3.8. ∎
Proof of Theorem 1.2.
By the assumptions, $A$ is a $\operatorname{GL}_{2}$-type abelian variety. Since the CM case is covered by e.g. [Fit24, Remark 12] and the case when $q=2$ is covered by [Saw16], we will thus assume that $A$ is not a CM abelian variety and $A$ is of dimension $q\geq 3$.
Let $F^{\prime}/F$ be a finite extension such that $\operatorname{End}_{F^{\prime}}(A)\otimes\mathbb{Q}=\operatorname{End}_{\overline{F}}(A)\otimes\mathbb{Q}$. Since $\dim A=q$ is a prime and $A$ is not a CM abelian variety, if $A/F^{\prime}$ is not simple, then by [Wu20, pp. 908–909 and Proposition 1.5], it must be isogenous over $F^{\prime}$ to either $B^{q}$ for some non-CM elliptic curve $B$ over $F^{\prime}$ or $B_{1}^{r_{1}}\times B_{2}^{r_{2}}$ for some CM abelian varieties $B_{1},B_{2}$ over $F^{\prime}$ with $r_{1}\dim B_{1}=r_{2}\dim B_{2}=q/2$. As $q$ is an odd prime, the latter case is not possible. In the former case, the set of ordinary primes of $B$ over $F^{\prime}$ has density $1$, so the set of ordinary primes of $A$ over $F$ has density at least $\frac{1}{[F^{\prime}:F]}$.
Now, suppose that $A/F^{\prime}$ is simple. Note that $q$ is an odd prime, $A$ is not a CM abelian variety, and $K\hookrightarrow\operatorname{End}_{F^{\prime}}(A)\otimes\mathbb{Q}$. By [Wu20, Proposition 1.7], we must have that $K=\operatorname{End}_{F^{\prime}}(A)\otimes\mathbb{Q}$ is a totally real field. In particular,
$K=\operatorname{End}_{F^{\prime}}(A)\otimes\mathbb{Q}=\operatorname{End}_{\overline{F}}(A)\otimes\mathbb{Q}.$
This implies that $A/\overline{F}$ must be simple as $\operatorname{End}_{\overline{F}}(A)\otimes\mathbb{Q}$ is a field. In this way, $A$ is absolutely simple, $\operatorname{End}_{F}(A)\otimes\mathbb{Q}$ contains a number field $K$ of degree $q=\dim A$, and $A/\overline{F}$ is not of type IV (in fact, $A$ is of type I), so $A$ satisfies the assumptions in Theorem 1.3. As $F$ is a quadratic field and $K$ is a number field whose degree is an odd prime, it follows from Corollary 2.10 that the condition (P1) in Theorem 1.3 is satisfied. Hence, the set of ordinary primes of $A$ over $F$ has positive density by Theorem 1.3. ∎
6.2. Application to modular forms
We will now prove Theorem 1.5 and we adopt all the notations from Theorem 1.5 and Section 3.2.
First, note that if $f$ is CM, then the set of ordinary primes of $f$ has positive density by Corollary A.2. We will thus assume that $f$ is non-CM from now on. By the discussions in Section 3.2, the family of $\ell$-adic Galois representations $\{\rho_{f,\ell}:\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})\rightarrow\operatorname{GL}_{2}(K_{f}\otimes\mathbb{Q}_{\ell})\}_{\ell\in\Sigma_{\mathbb{Q}}}$ satisfies that
  1. (1)
    each $\rho_{f,\ell}$ is unramified at rational primes $p$ with $p\nmid N\ell$;
  2. (2)
    for every rational primes $p\nmid N\ell$, $\operatorname{det}(XI-\rho_{f,\ell}(\operatorname{Frob}_{p}))=X^{2}-(a_{p}(f)\otimes 1)X+(\epsilon(p)p\otimes 1)$, which is independent of $\ell$;
  3. (3)
    $|\iota(a_{p}(f))|_{\mathbb{C}}\leq 2\sqrt{p}$ for all $p\nmid N$ and all embeddings $\iota:K_{f}\hookrightarrow\mathbb{C}$;
  4. (4)
    the set $\{p\nmid N:a_{p}(f)=0\}$ of rational primes has density 00;
  5. (5)
    there exist two finite extensions of fields $F^{\prime}_{f}/\mathbb{Q}$ and $K_{f}/K^{\prime}_{f}$ such that for $\ell$ sufficiently large, we have, up to conjugation, that
    $\rho_{f,\ell}(\operatorname{Gal}(\overline{\mathbb{Q}}/F^{\prime}_{f}))=\{M\in\operatorname{GL}_{2}(\mathcal{O}_{K^{\prime}_{f}}\otimes\mathbb{Z}_{\ell}):\operatorname{det}M\in\mathbb{Z}_{\ell}^{\times}\}.$
Here the only property that is not recalled in Section 3.2 is (4), which follows from [Ser81, Section 7.2].
From these properties, we can already deduce the second part of Theorem 1.5 by applying Theorem 1.6 (with $T$ being the set of rational primes that split completely in $K_{f}$) or Theorem 1.7. The first part of Theorem 1.5, i.e., when $[K_{f}:\mathbb{Q}]=q$ or $2q$ for some prime $q$, now follows from Proposition 2.9 and the second part.
Appendix A Ordinary primes of CM modular forms of weight $2$
In this appendix, we will prove that CM modular forms of weight $2$ have positive density ordinary primes. In fact, we will prove a result purely for certain Galois representations, from which the result for CM modular forms of weight $2$ follows. The result should already be known to experts, and the approach we take here is essentially the same as [Fit24, Remark 12].
Proposition A.1.
Let $F$ and $K$ be two number fields, let $\lambda$ be a fixed prime of $K$, and let $\epsilon$ be a Hecke character of $F$ of finite order with values in $K^{\times}$. Let $\rho:\operatorname{Gal}(\overline{F}/F)\rightarrow\operatorname{GL}_{2}(K_{\lambda})$ be a continuous Galois representation such that for all primes $v$ of $F$ outside a finite set $S$ of primes of $F$, $\rho$ is unramified at $v$ and
$\operatorname{det}(XI-\rho(\operatorname{Frob}_{v}))=X^{2}-\operatorname{tr}\rho(\operatorname{Frob}_{v})X+\epsilon(v)\mathrm{N}(v)\in\mathcal{O}_{K}[X].$
Suppose that there exists a finite extension $F^{\prime}/F$ such that for all but finitely many primes $v^{\prime}$ of $F^{\prime}$, $\rho|_{\operatorname{Gal}({\overline{F}/F^{\prime}})}$ is unramified at $v^{\prime}$ and $\operatorname{det}(XI-\rho|_{\operatorname{Gal}({\overline{F}/F^{\prime}})}(\operatorname{Frob}_{v^{\prime}}))$ splits into linear factors over $K$. Then,
$S^{\mathrm{ord}}:=\{v\in\Sigma_{F}:v\notin S\text{ and }\mu\nmid(\operatorname{tr}\rho(\operatorname{Frob}_{v}))\text{ for all }\mu\in\Sigma_{K}\text{ with }\mu|\mathrm{N}(v)\}$
has density greater than or equal to $\frac{1}{[F^{\prime}:F]}$.
Proof.
We first extend the finite set $S$ to include primes of $F$ that ramify in $F^{\prime}/F$ or lie above rational primes that ramify in $K/\mathbb{Q}$. Now, let $S^{\prime}$ be a finite set of primes of $F^{\prime}$ that contains all the ramified primes of $\rho^{\prime}:=\rho|_{\operatorname{Gal}(\overline{F}/F^{\prime})}$ and all the primes that lie above primes in $S$.
Let $v^{\prime}\notin S^{\prime}$ be a prime of $F^{\prime}$ of degree $1$, let $v$ be the prime of $F$ lying below $v^{\prime}$, and let $p=\mathrm{N}(v^{\prime})=\mathrm{N}(v)$ be the rational prime lying below $v^{\prime}$. As $v^{\prime}$ is of degree $1$, the residue fields of $v^{\prime}$ and $v$ coincide, so there exists a Frobenius lift at $v^{\prime}$ in $\operatorname{Gal}(\overline{F}/F^{\prime})$ such that it is also a Frobenius lift at $v$ in $\operatorname{Gal}(\overline{F}/F)$. Hence,
$\operatorname{det}(XI-\rho(\operatorname{Frob}_{v^{\prime}}))=\operatorname{det}(XI-\rho(\operatorname{Frob}_{v}))=X^{2}-a_{v}X+\epsilon(v)p=(X-\alpha_{v})(X-\beta_{v})$
for $a_{v}\in K$ and $\alpha_{v},\beta_{v}\in K$ (by the assumptions). Suppose that $v\notin S^{\mathrm{ord}}$. Then, there exists a prime $\mu$ of $K$ with $\mu|p$ such that $\mu|(a_{v})$. We thus have
$v_{\mu}(a_{v})\geq 1\quad\text{ and }\quad v_{\mu}(\epsilon(v)p)\geq 1.$
These imply that $v_{\mu}(\alpha_{v})\geq 1$ and $v_{\mu}(\beta_{v})\geq 1$. As $\epsilon(v)$ is a unit,
$v_{\mu}(p)=v_{\mu}(\epsilon(v)p)=v_{\mu}(\alpha_{v})+v_{\mu}(\beta_{v})\geq 2,$
so $p$ ramifies in $K/\mathbb{Q}$. This contradicts that $v^{\prime}\notin S^{\prime}$ by the definition of $S^{\prime}$. Thus, $v\in S^{\mathrm{ord}}$.
To summarize, if a prime $v$ of $F$ lies below a degree $1$ prime $v^{\prime}\notin S^{\prime}$ of $F^{\prime}$, then $v\in S^{\mathrm{ord}}$, i.e.,
$S_{0}:=\{v\in\Sigma_{F}:v^{\prime}|v\text{ for some degree $1$ prime $v^{\prime}\notin S^{\prime}$ of $F^{\prime}$}\}\subseteq S^{\mathrm{ord}},$
so it suffices to show that $S_{0}$ has density greater than or equal to $\frac{1}{[F^{\prime}:F]}$. Note that each $v\in S_{0}$ may split into at most $[F^{\prime}:F]$ primes of $F^{\prime}$ of degree $1$, so
$[F^{\prime}:F]\cdot\#\{v\in S_{0}:\mathrm{N}(v)\leq X\}\geq\#\{\text{degree $1$ prime $v^{\prime}$ of $F^{\prime}$}:\mathrm{N}(v^{\prime})\leq X\}.$
On the other hand, by Landau’s prime ideal theorem,
#$\#\{v\in\Sigma_{F}:\mathrm{N}(v)\leq X\}\sim\frac{X}{\log X}\sim\#\{v^{\prime}\in\Sigma_{F^{\prime}}:\mathrm{N}(v^{\prime})\leq X\}.$
Hence,
$\mathrm{Density}(S_{0})=\lim_{X\to\infty}\frac{\#\{v\in S_{0}:\mathrm{N}(v)\leq X\}}{\#\{v\in\Sigma_{F}:\mathrm{N}(v)\leq X\}}$ $\geq\frac{1}{[F^{\prime}:F]}\lim_{X\to\infty}\frac{\#\{\text{degree $1$ prime $v^{\prime}$ of $F^{\prime}$}:\mathrm{N}(v^{\prime})\leq X\}}{\#\{v\in\Sigma_{F}:\mathrm{N}(v)\leq X\}}$
$=\frac{1}{[F^{\prime}:F]}\lim_{X\to\infty}\frac{\#\{\text{degree $1$ prime $v^{\prime}$ of $F^{\prime}$}:\mathrm{N}(v^{\prime})\leq X\}}{\#\{v^{\prime}\in\Sigma_{F^{\prime}}:\mathrm{N}(v^{\prime})\leq X\}}$
$=\frac{1}{[F^{\prime}:F]}.$
We are now ready to prove the result for CM modular forms of weight $2$. We will take all the facts on CM modular forms for granted and refer readers to [Rib77, Section 3 and 4] for references.
Corollary A.2.
The set of ordinary primes of any CM newform of weight $2$ has density $1/2$.
Proof.
Let $f$ be a CM newform of weight $2$. Let $K_{f}$ be the coefficient field of $f$, let $\lambda$ be a prime of $K_{f}$, and let $\rho_{f,\lambda}:\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})\rightarrow\operatorname{GL}_{2}(K_{f,\lambda})$ be the $\lambda$-adic representation attached to $f$.
First, since there exists an imaginary quadratic field $F^{\prime}$ such that $a_{p}(f)=0$ for all but finite many primes $p$ that are inert in $F^{\prime}$, it follows that the set of ordinary primes of $f$ has density at most $1/2$. For the lower bound, note that there exists an algebraic Hecke character $\psi$ of $F^{\prime}$ such that $\rho_{f,\lambda}\simeq\mathrm{Ind}_{F^{\prime}/\mathbb{Q}}\psi$ (after extending the scalar). In particular, there exists a number field $K$ (where the image of $\psi$ lies) such that for all but finite many primes $v^{\prime}$ of $F^{\prime}$, $\rho^{\prime}:=\rho_{f,\lambda}|_{\operatorname{Gal}(\overline{\mathbb{Q}}/F^{\prime})}$ is unramified at $v^{\prime}$ and $\operatorname{det}(XI-\rho^{\prime}(\operatorname{Frob}_{v^{\prime}}))$ splits into linear factor over $K$. As $f$ is of weight $2$, the data $(\rho_{f,\lambda},F^{\prime}/\mathbb{Q},K)$ satisfies the condition in Proposition A.1, so it follows that the set of ordinary primes of $f$ has density at least $1/2$. The result then follows. ∎